### v2

# Infrequent Resolving Algorithm for Online Linear Programming

Abstract
Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management and advertising. Existing OLP algorithms fall into two categories: LP-based algorithms and LP-free algorithms. The former one typically guarantees better performance, even offering a constant regret, but requires solving a large number of LPs, which could be computationally expensive. In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound. In this work, we bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$. Moreover, when we are allowed to solve LPs only $M$ times, we propose an algorithm that can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret. Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Numerical experiments are conducted to demonstrate the efficiency of the proposed algorithms.
Keywords: online linear programming; network revenue management; resolving.
1 Introduction
Online linear programming (OLP) is a classical problem in online decision making. In this problem, a decision maker manages multiple types of resources, such as airplane seats or hotel rooms, with limited inventory. Customers arrive sequentially, each requesting a combination of resources (e.g., multi-leg flights or multi-day room reservations) and offering a bid price. Upon observing an arriving customer’s request and bid price, the decision maker needs to make an irrevocable decision to accept or reject the request in order to maximize the total expected revenue without violating resource constraints. OLP has applications in various industries, including online auction (Buchbinder et al. 2007), advertisement (Mehta et al. 2005), covering and packing (Buchbinder and Naor 2009b), e-commerce fulfillment (Jasin and Sinha 2015), and network revenue management (Jasin 2015), among many others. To facilitate our discussion, we first provide the formal formulation of the OLP problem.
Consider a decision maker with $m$ types of resources facing sequentially arriving customers over $T$ discrete time periods. We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{++}^{m}$ is given and denotes the vector of average inventories per period. Each customer arriving at period $t$ is characterized by a reward $r_{t}\in\mathbb{R}_{+}$ and a consumption vector $\boldsymbol{A}_{t}\in\mathbb{R}_{+}^{m}$. The time periods are indexed forward, starting from period $t=1$ and ending at period $t=T$. In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) [^1] Our results can be easily extended to the case with multiple actions. without knowing future information. Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) linear program:
$\max_{\boldsymbol{x}}\quad$ $\sum_{t=1}^{T}r_{t}x_{t}$
s.t. $\sum_{t=1}^{T}\boldsymbol{A}_{t}x_{t}\leq T\boldsymbol{\rho},$ (1)
$x_{t}\in\{0,1\},\quad\forall\;t.$
In the corresponding online setting, at each period $t$, the coefficients $(r_{t},\boldsymbol{A}_{t})$ in (1) are revealed, and the decision maker needs to determine the decision variable $x_{t}$ without knowing future information $\{(r_{\ell},\boldsymbol{A}_{\ell}):\ell=t+1,t+2,\dots,T\}$. The goal is to find a non-anticipative policy to maximize the expected revenue over the entire time horizon. Such a problem is commonly referred to as online linear programming (OLP).
In this paper, we assume the customers’ features $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$ with $n\in\mathbb{Z}_{+}$ supports. Such an assumption is widely adopted in the literature (e.g., Jasin 2015, Xie et al. 2023 and Chen et al. 2024). With this assumption, the OLP problem can be simplified as follows:
There are $n$ types of arrivals with corresponding coefficient groups $\{(r_{j},\boldsymbol{A}_{j})\}_{j=1}^{n}$, and each arrival’s type is drawn according to the distribution $\{p_{j}\}_{j=1}^{n}$ (with $\sum_{j=1}^{n}p_{j}=1$). Slightly abusing the notations, we define $\boldsymbol{r}\in\mathbb{R}_{+}^{n}$ and $\boldsymbol{A}\in\mathbb{R}_{+}^{m\times n}$ as the reward vector and the consumption matrix for all types of customers, respectively. Let $j^{t}$ denote the random type of the $t$-th arrival. The decision maker’s problem is to select a non-anticipative policy $\mu$, to maximize the total expected rewards while satisfying the inventory constraints. The OLP problem can be written as follows:
$\begin{split}V^{*}(T):=\max_{\mu\in\Pi}\quad&\mathbb{E}\left[\sum_{t=1}^{T}\sum_{j=1}^{n}r_{j}x^{t}_{\mu,j}\right]\\ \text{s.t.}\quad&\sum_{t=1}^{T}\sum_{j=1}^{n}\boldsymbol{A}_{j}x^{t}_{\mu,j}\leq T\boldsymbol{\rho},\ (a.s.),\\ &x^{t}_{\mu,j}\leq Y_{j}^{t},\ \forall j,t,\ (a.s.),\\ &x^{t}_{\mu,j}\in\{0,1\},\ \forall j,t,\end{split}$ (2)
where $\Pi$ is the set of all non-anticipative policies, $x^{t}_{\mu,j}$ denotes whether the policy $\mu$ accepts a type-$j$ customer at time $t$, and $Y_{j}^{t}\in\{0,1\}$ is the random indicator whether a type-$j$ customer arrives at time $t$. If we relax the integer constraints in (2) and replace all random variables with their expectations, then the model is referred to as the “fluid model”.
Solving the optimal policy $\mu^{*}$ for (2) is computationally intractable even if the underlying distribution $\mathcal{P}$ is known. To address such challenges, researchers have proposed various heuristic policies with performance guarantees. A widely adopted performance measure is regret, which quantifies the optimality gap of a policy. Specifically, let $V^{\mu}(T)$ denote the expected revenue accumulated over the time horizon $T$ under a given policy $\mu$. We define the regret formally as follows:
$\text{Reg}^{\mu}(T)=V^{*}(T)-V^{\mu}(T)=V^{*}(T)-\mathbb{E}\left[\sum_{t=1}^{T}\sum_{j=1}^{n}r_{j}x_{\mu,j}^{t}\right].$ (3)
In an asymptotic regime where $T$ scales up, a policy $\mu$ is asymptotically optimal if $\text{Reg}^{\mu}(T)=o(T)$, implying the competitive ratio $V^{\mu}(T)/V^{*}(T)$ converges to one as the time horizon $T$ goes to infinity. Furthermore, a lower-order regret typically implies a better performance, and designing a policy with constant regret, i.e., $\text{Reg}^{\mu}(T)=\mathcal{O}(1)$, is highly desirable.
Broadly speaking, OLP algorithms in prior studies can be categorized into two types: LP-based algorithms and LP-free algorithms. LP-based algorithms typically make decisions based on the optimal primal/dual solution to the fluid LP in each period, offering strong performance guarantees but requiring solving a large number of LPs (e.g., Li and Ye 2022, Xie et al. 2023, Chen et al. 2024, etc.). Although commercial solvers can efficiently solve LPs, frequent LP resolving remains computationally expensive especially for large-scale or time-sensitive problems. In response, researchers recently propose LP-free algorithms that typically use gradient descent methods to derive an approximate solution converging to the optimal dual solution to the fluid LP, and makes decisions based on the approximate solution. Thus, such algorithms rely only on first-order computations and avoid solving any LP (e.g., Li et al. 2020, Gao et al. 2024, Ma et al. 2024, etc.). However, these LP-free approaches usually lead to weaker performance bound than the LP-based algorithms.
In this paper, we aim to strike a balance between the performance guarantee of LP-based algorithms and the computational efficiency of LP-free algorithms. Specifically, we propose an algorithm which achieves $\mathcal{O}(1)$ regret for the OLP problem while requiring solving LPs only $\mathcal{O}(\log\log T)$ times. In our algorithm, we concentrate the resolving (of an approximate LP) periods 1) on the beginning of the time horizon and 2) toward the end of the time horizon. The first set of resolvings aim to update the policy when the data is scarce so that correcting learning errors is important, while the latter set of resolvings aim to update the policy when the inventory is running out. Between two resolvings, we use the solution of an approximate LP to guide the allocation. Moreover, when we are allowed to solve LPs for up to $M$ times, we propose an algorithm that can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret. In addition, we show that the proposed algorithms can be easily adapted to the settings where the arrival probabilities of each type of request are known, which also lead to near-optimal regrets in those settings. In this case, a modified algorithm can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times. Therefore, our results demonstrate that one can achieve near optimal regrets only with a few resolvings, and depict the precise tradeoff between the frequency of resolving and the performance of the algorithm.
We would like to highlight the technical contribution of this work. In the literature of LP-based algorithms for solving OLP, the algorithm usually solves an updated fluid model in each time period (e.g., Vera and Banerjee 2021, Xie et al. 2023, Chen et al. 2024). This approach allows bridging their policies with the optimal hindsight policy through two LPs: the fluid model and the hindsight benchmark. In contrast, our approach solves the updated fluid model only at a few selected periods ($\mathcal{O}(\log\log T)$ periods), and hence cannot access the optimal solution of the updated fluid model in most periods, posing a challenge for the analysis. In order to overcome this technical challenge, we approximate the optimal solution of the updated fluid model based on the latest obtained fluid solution and some first-order computations. While this solution is sub-optimal for the updated fluid model, we prove that it is optimal to a surrogate LP with high probability. Therefore, we can bridge our policy with the optimal hindsight policy through the surrogate LP and the hindsight LP, from which we can obtain the desired result.
The remainder of this paper is organized as follows. In the rest of this section, we review literature related to our work. In Section 2, we propose the main algorithm and prove the regret bounds under the infrequent resolving schedule and the finite-resolving schedule. In Section 3, we study the case with known arrival probabilities, which is referred to as the known-probability case. In Section 4, we compare our policy with several existing policies and provide additional insights through numerical experiments. We conclude the paper in Section 5. All proofs are relegated to the Appendix.
1.1 Literature Review
Online decision making has a rich history within operations research and theoretical computer science, and remains a vibrant and flourishing area. Academic studies in this field typically focus on designing online algorithms that make real-time decisions based on limited information and adapt their strategies as new data becomes available. For a comprehensive review, we refer readers to Borodin and El-Yaniv (2005), Buchbinder and Naor (2009a) and Hazan (2016). Online linear programming (OLP) problem is a classical problem in online decision making. In the OLP problem, in order to maximize the expected revenue under resource constraints, the decision maker needs to dynamically make irrevocable decisions to accept or reject the arriving customers’ requests. There are two streams of research categorized by whether the arrival probabilities is known at the beginning. In the following, we review the literature of these two streams separately.
Unknown Distribution. We start with the stream assuming unknown arrival probabilities. Our work closely relates to studies under the random input assumption, where the coefficients $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d. from an unknown distribution $\mathcal{P}$. We focus on the asymptotic order of the regret as $T$ increases. Bray (2024) shows that the regret is at least $\mathcal{O}(\log T)$ when the demand distribution has infinite supports. In this paper, we focus on the case with a finite-support demand distribution. We first review the so called LP-based algorithms which require solving many LPs but guarantee good performance. For example, under certain non-degeneracy assumptions, Jasin (2015) proposes an $\mathcal{O}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times. Subsequently, Chen et al. (2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise. Under the non-degeneracy assumption and some distribution assumptions, Li and Ye (2022) provide an $\mathcal{O}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times. Then, under similar assumptions, Ma et al. (2024) provide an $\mathcal{O}(\log T)$-regret algorithm requiring solving LPs $\mathcal{O}(T)$ times. Recently, Xie et al. (2023) remove the non-degeneracy assumption and provide an OLP algorithm with $\mathcal{O}(1)$ regret but still requiring solving LPs $T$ times. Moreover, there are studies on infrequent resolving algorithms under additional assumptions. Given a complete sample path (i.e., $T$ samples), Freund and Banerjee (2019) propose an algorithm with lazy resolving (i.e., $\mathcal{O}(\log\log T)$ expected LP solving times) and prove a constant regret. Given a starting point with an accuracy of $\mathcal{O}(1/T)$, Ma et al. (2024) propose an infrequent resolving algorithm that can achieve $\mathcal{O}(\log T)$ regret under the non-degeneracy assumption by solving LPs $O(\log T)$ times. Both assumptions are stringent when $T$ is large. In this work, without such assumptions, we propose an algorithm that can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times.
Table 1: Comparison among OLP Algorithms (i.e., Unknown-Probability Case).
Paper
Regret
(Non-degenerate Case)
Regret
(Degenerate Case)
# of Resolvings
Jasin (2015) $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log T)$
Li and Ye (2022) $\mathcal{O}(\log T\log\log T)$ $\mathcal{O}(T)$
Chen et al. (2024) $\mathcal{O}(1)$ $\mathcal{O}(\sqrt{T}\log T)$ $\mathcal{O}(T)$
Xie et al. (2023) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Ma et al. (2024) $\mathcal{O}(\log T)$ $\mathcal{O}(T)$
Freund and Banerjee (2019)
(Requiring $T$ samples )
$\mathcal{O}(1)$ $\mathcal{O}(1)$
$\mathcal{O}(\log\log T)$
(In expectation )
Ma et al. (2024)
(Requiring $\mathcal{O}(1/T)$-accuracy starting point )
$\mathcal{O}(\log T)$ $\mathcal{O}(\log T)$
Li et al. (2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00
Gao et al. (2024) $\mathcal{O}(T^{1/3})$ 00
Ma et al. (2024) $\mathcal{O}(\log^{2}T)$ 00
This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $M$
Although LP-based algorithms have good performance guarantees, frequent LP solving can be computational expensive for large-scale or time-sensitive problems. To address these computational challenges, recent studies have developed LP-free algorithms that never solve full LPs. Most of these studies analyze the regret in absence of budget constraints and derive bounds for both the regret and the constraint violation, i.e., the cumulative over-budget consumption. [^2] As Li et al. (2020) note, although an LP-free algorithm can be modified to a feasible policy satisfying the budget constraints, deriving a regret bound for such a modified policy remains an open question. The pioneering work Li et al. (2020) adopts a stochastic sub-gradient descent method to learn the optimal dual prices and provides an $\mathcal{O}(\sqrt{T})$-regret LP-free algorithm with $\mathcal{O}(\sqrt{T})$ constraint violation. This algorithm only requires first-order computations and never solves any full LP. Subsequently, researchers have studied variants of this algorithm and derived tighter bounds under the non-degeneracy assumption and the distribution assumptions in Li and Ye (2022). For example, Gao et al. (2024) propose variants that improve the bounds for both regret and the constraint violation to $\mathcal{O}(T^{1/3})$. Similarly, Ma et al. (2024) propose an LP-free algorithm which guarantees an $\mathcal{O}(\log^{2}T)$ regret without constraint violation. In this work, we achieve a balance between computational efficiency and algorithm performance by proposing an algorithm that achieves a constant regret bound by solving LPs $\mathcal{O}(\log\log T)$ times. A summary of the existing results and our result is presented in Table 1.
Another stream of research studies the OLP problem under the random permutation assumption. Here, the set $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$ is adversarially chosen, but the arrival order is uniformly distributed over all the permutations. Researchers aim to derive necessary conditions for the existence of a $(1-\epsilon)$-competitive algorithm, see, e.g., Agrawal et al. (2014), Molinaro and Ravi (2014), Gupta and Molinaro (2014), Kesselheim et al. (2014). Since the concentration bound under the random permutation assumption is weaker, the lower bound in Agrawal et al. (2014) implies that the asymptotic order of the regret is $\Omega(\sqrt{T})$. We also highlight that Agrawal et al. (2014), Molinaro and Ravi (2014) and Gupta and Molinaro (2014) solve LPs $\mathcal{O}(\log T)$ times and their resolving periods concentrate at the beginning, sharing similarities with our resolving schedule.
Known distribution. We now review the stream assuming the knowledge of the arrival probabilities at the beginning, which is widely studied in the network revenue management (NRM) problem. Originating from the airline industry, NRM has garnered significant attention from both academia and industry. There are two mainstreams of research, price-based NRM and quantity-based NRM. The former one studies the dynamic pricing problem under resource constraints (see Gallego and Van Ryzin 1994); the latter one studies the dynamic resource allocation problem (see Talluri and Van Ryzin 1998). Our work is closely related to the quantity-based NRM, in which most algorithms are LP-based. For example, Talluri and Van Ryzin (1998) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the bid-price control (BPC) policy, which uses the optimal dual variable to the fluid model as the values of resources and accepts a request if and only if the offered price is larger than the total value of resources. Cooper (2002) proves an $\mathcal{O}(\sqrt{T})$ regret bound for the booking-limit control (BLC) policy, which assigns quotas to each request according to the fluid model and accepts a request until the corresponding quota is depleted. Reiman and Wang (2008) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the probabilistic allocation control (PAC) policy, which probabilistically accepts a request according to the ratio of the fluid model solution to the expected demand.
Table 2: Comparison among Algorithms for Known-Probability Case.
Paper
Regret
(Non-degenerate)
Regret
# of Resolvings
Jasin and Kumar (2012) $\mathcal{O}(1)$ [^3] Bumpensanti and Wang (2020) prove that the regret of the algorithm in Jasin and Kumar (2012) is $\Theta(\sqrt{T})$. $\mathcal{O}(\log T)$
Bumpensanti and Wang (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
Arlotto and Gurvich (2019)
(Single constraint )
$\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Vera and Banerjee (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Vera et al. (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
Sun et al. (2020) $\mathcal{O}(T^{3/8}(\log T)^{5/4})$ 0
Reiman and Wang (2008) $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ 2
Jasin and Kumar (2012) $\mathcal{O}(T^{\frac{1}{2^{M}}})$ $M$
Bumpensanti and Wang (2020) $\mathcal{O}(T^{5/12})$ $\mathcal{O}(T^{5/12})$ 2
Gupta (2024) $\mathcal{O}(1)$ 1
This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $M$
Observing the potential to resolve the fluid model to reduce the regret, many works consider algorithms with resolving. For example, Cooper (2002) provides an example where frequent resolving may hurt the revenue. Then, considering the asymptotic regime, Reiman and Wang (2008) prove that a single resolving can reduce the regret of the PAC policy from $\Theta(\sqrt{T})$ to $o(\sqrt{T})$. Jasin and Kumar (2012) and Jasin and Kumar (2013) show that the PAC policy can significantly benefit from resolving, reducing its regret from $\Theta(\sqrt{T})$ to $\mathcal{O}(1)$, while neither BPC nor BLC can benefit. Specifically, under the non-degeneracy assumption, the PAC policy with periodic ($\mathcal{O}(T)$ times) or midpoint ($\mathcal{O}(\log T)$ times) resolving can achieve $\mathcal{O}(1)$ regret. Without the non-degeneracy assumption, Bumpensanti and Wang (2020) show that the above regret in general is $\Omega(\sqrt{T})$, and provide a modified PAC policy with infrequent (i.e., $\mathcal{O}(\log\log T)$ times) resolving that guarantees $\mathcal{O}(1)$ regret. Recently, Arlotto and Gurvich (2019) study an alternative interpretation of the fluid solution, which accepts the request if and only if the ratio of the primal solution to the expected demand is no less than $1/2$. They prove an $\mathcal{O}(1)$ regret bound for the multi-secretary problem. In our work, we refer to such policy as the “argmax policy” because it takes the action (accept/reject) with the larger value in the primal solution. Vera and Banerjee (2021) and Vera et al. (2021) generalize this idea to the multi-constraint problem, but their policy requires solving LPs in every period. In our work, we also use the argmax policy as the base policy. We show that such a policy can guarantee $\mathcal{O}(1)$ regret by solving only $\mathcal{O}(\log\log T)$ LPs. In the LP-free regime, Sun et al. (2020) propose an LP-free algorithm with $\mathcal{O}(T^{3/8}(\log T)^{5/4})$ regret.
Finally, we would like to mention some results on the case when the decision maker is only allowed to solve a finite number (i.e., $M$) of LPs under the NRM model. For this problem, Reiman and Wang (2008) show that the regret can be reduced to $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ if we can solve LPs twice. However, the proof cannot be directly extended to the multiple-resolving case. Under the non-degeneracy assumption, Jasin and Kumar (2012) prove that solving LP at $M$ periods can induce an $\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret bound. Due to the restricted choice of resolving times, Bumpensanti and Wang (2020) prove that the regret bound of their policy is $\mathcal{O}(T^{5/12})$ given $M=2$, which is worse than previous results. In addition, Gupta (2024) recently proposes an algorithm which only solves LP once at the beginning and then greedily makes decisions to minimize a sum-of-squares function in each period, whose regret is proved to be $\mathcal{O}(1)$ under the non-degeneracy assumption. Compared to the above literature, our results can meet the single-resolving result of Reiman and Wang (2008) and extend it to the finite-resolving case. We derive similar regret bounds in Jasin and Kumar (2012) but allow for degenerate cases. A summary of the existing results and our result is presented in Table 2 (where $\epsilon$ can be any positive constant).
2 Main Results
In this section, we propose an algorithm to solve problem (2) with constant regret. For the ease of notation, we define the following LP parameterized by inventory $\boldsymbol{b}$ and demand $\boldsymbol{d}$:
$\begin{split}\phi(\boldsymbol{b},\boldsymbol{d}):=\max_{\boldsymbol{y}\geq\boldsymbol{0}}\quad&\boldsymbol{r}^{\mathsf{T}}\boldsymbol{y}\\ \text{s.t.}\quad&\boldsymbol{A}\boldsymbol{y}\leq\boldsymbol{b}\\ &\boldsymbol{y}\leq\boldsymbol{d}.\end{split}$ (4)
At time $t$, suppose the remaining inventory is $\boldsymbol{b}^{t}$ and the realized demand from period $\ell=1$ to period $\ell=t-1$ are $\{Y^{\ell}_{j}:\ell=1,2,\dots,t-1,j=1,2,\dots,n\}$. We refer to $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ as the “fluid model" at time $t$, where $\hat{p}_{j}^{t}=\left(\sum_{\ell=1}^{t-1}Y_{j}^{\ell}\right)/(t-1)$ is the empirical estimation of arrival probability $p_{j}$ at time $t$. We also let $\hat{\boldsymbol{p}}^{1}=\boldsymbol{0}$. The fluid model replaces all uncertainties with their expectations. The decision variable $y_{j}$ represents the expected number of accepted type-$j$ customer. The first constraint ensures that the total resource consumption does not exceed the remaining inventory, and the second constraint ensures that the number of accepted customers does not exceed the demand. For the second constraint, since the arrival probabilities $p_{j}$ are unknown, we use the empirical estimation $\hat{\boldsymbol{p}}^{t}$, approximating the future demand as $(T-t+1)\hat{\boldsymbol{p}}^{t}$.
2.1 Argmax with Infrequent Resolving (AIR) Policy
We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy. The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
Algorithm 1 Argmax with Infrequent Resolving (AIR) Policy
Input: Time set $\mathcal{T}=\{T_{1},T_{2},T_{3},\dots,T_{|\mathcal{T}|}\}$.
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$, $\boldsymbol{N}^{1}\leftarrow\boldsymbol{0}$, $\boldsymbol{u}^{1}\leftarrow\boldsymbol{0}$ and $\boldsymbol{d}^{1}\leftarrow\boldsymbol{0}$.
for $t=1,2,3,\dots,T$ do
     if $t\in\mathcal{T}$ then $\triangleright$ Infrequent resolving
         Update the empirical estimations $\hat{p}_{j}^{t}\leftarrow N_{j}^{t}/(t-1)$ for each $j$.
         Solve the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ and obtain its optimal solution $\boldsymbol{y}^{*}$.
         Set $u^{t}_{j}\leftarrow y^{*}_{j}$ and $d^{t}_{j}\leftarrow\hat{p}^{t}_{j}(T-t+1)$ for all $j$.
     end if
     Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
         Accept the request.
         Set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$. $\triangleright$ Update the remaining capacity
         Set $\boldsymbol{u}^{t+1}\leftarrow\boldsymbol{u}^{t}-\boldsymbol{e}_{j}$. $\triangleright$ Approximate the optimal solution
     else
         Reject the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}$.
     end if
     Set $\boldsymbol{d}^{t+1}\leftarrow\boldsymbol{d}^{t}-\boldsymbol{e}_{j}$$\triangleright$ Approximate the future demand
end for
We now explain the intuition of Algorithm 1. The algorithm begins with a predetermined time set $\mathcal{T}$ specifying the time points where resolving is needed. At each time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ as an approximation to the future demand and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$. The optimal solution $\boldsymbol{y}^{*}$ of the fluid LP provides an approximation to the optimal policy. At each time period $t\notin\mathcal{T}$, the decision maker updates an approximation $\boldsymbol{u}^{t}$ to the optimal solution in a “greedy” fashion as described in the algorithm. We adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera and Banerjee (2021) to make accept/reject decisions. At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand. Intuitively, $d_{j}^{t}-u_{t}^{j}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected. In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
In the following, we specify a resolving schedule $\mathcal{T}$ with $|\mathcal{T}|=\mathcal{O}(\log\log T)$, and then prove the constant regret bound.
2.2 Resolving Schedule
To achieve the constant regret, we introduce the time set $\mathcal{T}=\mathcal{T}_{L}\cup\mathcal{T}_{A}$. Specifically, the first subset is called the “learning” time set, specified as
$\mathcal{T}_{L}=\left\{\left\lceil T^{\alpha^{K_{L}}}\right\rceil,\dots,\left\lceil T^{\alpha^{3}}\right\rceil,\left\lceil T^{\alpha^{2}}\right\rceil,\left\lceil T^{\alpha}\right\rceil\right\}\bigcup\left\{\left\lceil\frac{T}{2}\right\rceil\right\}$
with $\alpha\in(0,1)$ and $K_{L}=\lceil\log_{\frac{1}{\alpha}}\log_{3}T\rceil$. The second subset is called the “approximation” time set, and is specified as
$\mathcal{T}_{A}=\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\left\lceil T-T^{\beta^{3}}\right\rceil,\dots,\left\lceil T-T^{\beta^{K_{A}}}\right\rceil\right\}$
with $\beta\in(\frac{1}{2},1)$ and $K_{A}=\lceil\log_{\frac{1}{\beta}}\log_{3}T\rceil$. To facilitate understanding, we illustrate the resolving times in Figure 1.
[…]
Figure 1: Illustration of Resolving Time Set $\mathcal{T}=\mathcal{T}_{L}\cup\mathcal{T}_{A}$.
As Figure 1 shows, the learning time set $\mathcal{T}_{L}$ is concentrated at the beginning of the time horizon. This is similar to the approach in Agrawal et al. (2014) and Gupta and Molinaro (2014), but our method requires solving fewer LPs, i.e., $\mathcal{O}(\log\log T)$ times compared to $\mathcal{O}(\log T)$ times in Agrawal et al. (2014) and Gupta and Molinaro (2014). Since empirical estimations fluctuate drastically at the beginning of the time horizon, the learning time set is designed to update the estimations promptly to avoid the accumulation of error. In contrast to the learning time set, the approximation time set $\mathcal{T}_{A}$ is concentrated at the end of the time horizon. It shares some similarities with the NRM literature, e.g., Jasin and Kumar (2012) and Bumpensanti and Wang (2020). Since the solution to $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ usually changes drastically when the inventory is scarce, it is crucial to have a good approximation of the optimal solution at the end of the time horizon. The approximation time set is thus designed to control the deviation of the approximated solution from the true optimal fluid solution. The total number of resolving is $|\mathcal{T}|=K_{L}+K_{A}+1=\mathcal{O}(\log\log T)$. Then, we present our main result in the following theorem.
Theorem 1 (Regret Bound).
Given the resolving schedule $\mathcal{T}$ with $\alpha\in(0,1)$ and $\beta\in(\frac{1}{2},1)$, the regret of the AIR policy is $\mathcal{O}(1)$.
According to Theorem 1, the AIR algorithm can achieve a constant regret by solving $\mathcal{O}(\log\log T)$ LPs. In the following, we present the proof idea of Theorem 1.
2.3 Proof of Regret Bound
In this section, we provide a sketch of the proof of Theorem 1. In our discussions, without loss of generality, we assume the time periods in $\mathcal{T}$ are indexed in an ascending way, i.e., $T_{1}<T_{2}<\cdots<T_{|\mathcal{T}|}$. We then present some properties of the time set $\mathcal{T}$. Since $\min_{t\in\mathcal{T}}t\leq 3$ and $\max_{t\in\mathcal{T}}t\geq T-3$, we will focus on the properties for $t\in[3,T-3]$, such that there exist $T_{k-1}$ and $T_{k}$ satisfying $t\in[T_{k-1},T_{k})$. Then, we have the following lemma depicting the relation between $t$ and $T_{k-1}$.
Lemma 1.
Given the resolving schedule $\mathcal{T}$ and any $t\in[T_{k-1},T_{k})$, we have $T_{k-1}\geq(t-1)^{\alpha}$ and $T-T_{k-1}\leq(T-t+1)^{1/\beta}$.
According to Lemma 1, we can bound any time $t$ by the latest LP solving time $T_{k-1}$. Next we present the main steps in the proof. The proof can be decomposed into four steps as follows.
Step 1: Regret decomposition. To bound the regret (3), we start with a benchmark serving as an upper bound to the optimal value $V^{*}(T)$. It is well known that the hindsight problem provides a better approximation to NRM problem compared to the fluid model. Specifically, given full arrival information $\boldsymbol{Z}_{t}$ after time $t$, the hindsight problem at time $t$ is defined as $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$ where $\boldsymbol{b}^{t}$ is the vector of the remaining inventory at time $t$ with $\boldsymbol{b}^{1}=T\boldsymbol{\rho}$ being the initial inventory, $Z_{j}^{t}=\sum_{\ell=t}^{T}Y_{j}^{\ell}$ is the number of future type-$j$ arrivals after period $t$, and $\boldsymbol{Z}^{t}$ is the corresponding vector. Moreover, we have $\phi(\boldsymbol{b}^{T+1},\boldsymbol{Z}^{T+1})=0$ because $\boldsymbol{Z}^{T+1}=\boldsymbol{0}$. Intuitively, the hindsight problem has access to future information and thus gains more rewards comparing to any non-anticipative online algorithm. The following result formalizes this intuition.
Lemma 2 (Upper Bound).
Given any $\boldsymbol{b}^{1}\geq\boldsymbol{0}$, we have $\mathbb{E}[\phi(\boldsymbol{b}^{1},\boldsymbol{Z}^{1})]\geq V^{*}(T).$
According to Lemma 2, it is sufficient to bound the regret by analyzing the difference between the hindsight problem and the performance of our policy. Let $\mathcal{A}$ denote the AIR policy, we have
$\begin{split}\text{Reg}^{\mathcal{A}}(T)&=V^{*}(T)-\mathbb{E}\left[\sum_{t=1}^{T}r_{j^{t}}x^{t}_{\mathcal{A},j^{t}}\right]\\ &\leq\mathbb{E}\left[\phi(\boldsymbol{b}^{1},\boldsymbol{Z}^{1})-\sum_{t=1}^{T}r_{j^{t}}x^{t}_{\mathcal{A},j^{t}}\right]\\ &=\mathbb{E}\left[\sum_{t=1}^{T}\left(\phi(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t})-\mathbb{E}\left[\phi(\boldsymbol{b}^{t+1}_{\mathcal{A}},\boldsymbol{Z}^{t+1})+r_{j^{t}}x^{t}_{\mathcal{A},j^{t}}\right]\right)\right]\\ &=\sum_{t=1}^{T}\mathbb{E}\left[\underbrace{\phi(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t})-\phi(\boldsymbol{b}^{t}_{\mathcal{A}}-x^{t}_{\mathcal{A},j^{t}}\boldsymbol{A}_{j^{t}},\boldsymbol{Z}^{t}-\boldsymbol{e}_{j^{t}})-r_{j^{t}}x^{t}_{\mathcal{A},j^{t}}}_{\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x^{t}_{\mathcal{A},j^{t}})}\right],\end{split}$ (5)
where $\boldsymbol{b}^{t}_{\mathcal{A}}=T\boldsymbol{\rho}-\sum_{\ell=1}^{t-1}x^{\ell}_{\mathcal{A},j^{\ell}}\boldsymbol{A}_{j^{\ell}}$ is the random inventory vector at time $t$ under the algorithm $\mathcal{A}$. Therefore, it is sufficient to bound $\mathbb{E}\left[\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})\right]$.
In the following, we introduce properties of the term $\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})$.
Proposition 1.
For the term $\Delta(\boldsymbol{b},\boldsymbol{Z},j,x)$, we have the following properties:
  1. (i)
    For any $(\boldsymbol{b},\boldsymbol{Z},j,x)$ with $\boldsymbol{b}\geq x\boldsymbol{A}_{j}$ and $\boldsymbol{Z}\geq\boldsymbol{e}_{j}$, we have $\Delta(\boldsymbol{b},\boldsymbol{Z},j,x)\leq r_{\phi}$, where $r_{\phi}$ is a constant independent of $T$.
  2. (ii)
    If there exists an optimal solution $\boldsymbol{y}^{*}$ to $\phi(\boldsymbol{b},\boldsymbol{Z})$ such that $y^{*}_{j}\geq 1$, then $\Delta(\boldsymbol{b},\boldsymbol{Z},j,1)=0$.
  3. (iii)
    If there exists an optimal solution $\boldsymbol{y}^{*}$ to $\phi(\boldsymbol{b},\boldsymbol{Z})$ such that $Z_{j}-y^{*}_{j}\geq 1$, then $\Delta(\boldsymbol{b},\boldsymbol{Z},j,0)=0$.
Proposition 1(i) states that the term $\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})$ is upper bounded by a constant $r_{\phi}$. More importantly, Proposition 1(ii) and (iii) provide conditions where the per-period optimality gap is zero. For example, if there exists an optimal solution $\boldsymbol{y}^{*}$ to $\phi(\boldsymbol{b},\boldsymbol{Z})$ such that $y_{j}^{*}\geq 1$, it is optimal for the clairvoyant to accept $y_{j}^{*}$ number of customers with type $j$. Since the reward is independent of time, the clairvoyant can accept the request either now or later without loss of optimality. Therefore, the per-period optimality gap of taking the acceptance action equals zero. Leveraging on this proposition, it holds that
$\text{Reg}^{\mathcal{A}}(T)\leq\sum_{t=1}^{T}\mathbb{E}\left[\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})\right]\leq\sum_{t=1}^{T}r_{\phi}\mathbb{P}\left(\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})>0\right),$ (6)
which reduces to bound the probability $\mathbb{P}\left(\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})>0\right)$. We will then show this probability is relatively small by arguing the sufficient conditions stated in Proposition 1(ii) and (iii) happen almost all the time. Notably, those sufficient conditions depend on the optimal solution to the hindsight problem at each period which is not available to the decision maker because resolving only happens at certain time points. Also, the conditions depend on the exact demand information which is not feasible to the decision maker. Therefore, we aim to show that under our resolving time schedule, those factors are well approximated. In the following, since the context is clear, we omit the dependence on policy $\mathcal{A}$ for $\boldsymbol{b}$ and $x$ in the notation.
Step 2: Bound demand approximation error. In this step, we prove that the approximated future arrivals $\boldsymbol{d}^{t}$ in Algorithm 1 is close to the true demand $\boldsymbol{Z}^{t}$.
Proposition 2 (Demand Approximation Error).
Given a time $t\in[T_{k-1},T_{k})$, we have
  1. (i)
    With probability larger than $1-\frac{2}{(t-1)^{2}}-\frac{2}{(T-t+1)^{2}}$, it holds that
    $|d_{j}^{t}-Z_{j}^{t}|\leq(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}+\sqrt{(T-T_{k-1}+1)\log(T-t+1)}.$
  2. (ii)
    With probability larger than $1-\frac{4}{(T-t+1)^{2}}$, it holds that
    $|d_{j}^{t}-Z_{j}^{t}|\leq(T-T_{k-1}+1)\sqrt{\frac{\log(T-t+1)}{T_{k-1}-1}}+\sqrt{(T-T_{k-1}+1)\log(T-t+1)}.$
According to Proposition 2, the demand estimation error is relatively small with high probability. Both bounds consist of two parts: The first term is due to the gap between the empirical probability $\hat{\boldsymbol{p}}^{t}$ and the underlying true probability $\boldsymbol{p}$; the second term is due to the deviation of the random future demand $\boldsymbol{Z}^{t}$ from the expected demand $(T-t+1)\hat{\boldsymbol{p}}^{t}$.
Step 3: Surrogate LP for $\boldsymbol{u}^{t}$. In the literature on argmax policies (e.g., Vera and Banerjee 2021 and Vera et al. 2021), the proposed policy makes decisions based on the optimal solution of the fluid problem at each period. This approach allows them to bridge the proposed policy with the optimal policy using two LPs: the fluid problem and the hindsight problem. However, we solve the fluid problem only at a few selected periods, and approximate the optimal solution by $\boldsymbol{u}^{t}$ in other periods. In this case, $\boldsymbol{u}^{t}$ is not an optimal solution to the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ for most time periods. Consequently, the linkage between the fluid model and the proposed policy breaks down, and the proof technique in the literature cannot directly apply to our proof. To overcome this challenge, we prove that $\boldsymbol{u}^{t}$ is a good approximation to the optimal solutions of the fluid problem by introducing a surrogate LP. In the following, we show that $\boldsymbol{u}^{t}$ is an optimal solution to the surrogate LP with high probability.
Proposition 3 (Surrogate LP).
Given the AIR policy with the resolving schedule $\mathcal{T}$ with $\alpha\in(0,1)$ and $\beta\in(\frac{1}{2},1)$, there exist constants $c_{1}$ and $c_{2}$ independent of $\;T$ such that when $t\in[c_{1},T-c_{2}]$, with probability larger than $1-\frac{n}{(T-t+1)^{2}}-\frac{n}{(\min\{T-t+1,t-1\})^{2}}$, we have
  1. (i)
    $\boldsymbol{d}^{t}\geq\frac{T-t+1}{2}\boldsymbol{p}\geq\boldsymbol{2}$.
  2. (ii)
    $\boldsymbol{u}^{t}$ is an optimal solution of the LP $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$.
As Proposition 3 shows, the approximated solution $\boldsymbol{u}^{t}$ is optimal to a surrogate LP, $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$, with high probability. Recall that we prove that the demand estimation $\boldsymbol{d}^{t}$ is not far away from the random demand $\boldsymbol{Z}^{t}$. Then, we will prove that the optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$ is not far away from the optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$. Thus, we can show that if $u_{j}^{t}\geq d_{j}^{t}/2$, then there exists an optimal solution $\boldsymbol{y}^{*}$ of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$ such that $y^{*}_{j}\geq 1$ with high probability, implying that $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j,1)=0$ according to Proposition 1. The logic for the case with $u_{j}^{t}<d_{j}^{t}/2$ is similar.
Step 4: Bound the probability $\mathbb{P}\left(\Delta(\boldsymbol{b}^{t}_{\mathcal{A}},\boldsymbol{Z}^{t},j^{t},x_{\mathcal{A},j^{t}}^{t})>0\right)$. We first define the “good event” at period $t\in[c_{1},T-c_{2}]$ to be that the conditions in both Propositions 2 and 3 hold for all $j$ in this period, and let the “bad event” denote the opposite. The probability of good event at period $t$ is at least $1-\frac{c_{3}}{(T-t+1)^{2}}-\frac{c_{4}}{(t-1)^{2}}$ where $c_{3}$ and $c_{4}$ are positive constants independent of $T$. In the following, we will prove that $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$ conditioned on the good event for most periods.
According to Proposition 1, in order to show $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$, it suffices to show that there exists an optimal solution $\boldsymbol{y}^{*}$ of $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$ such that $y^{*}_{j^{t}}\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j^{t}}^{t}-y^{*}_{j^{t}}\geq 1$ otherwise. As Proposition 3 shows, we can bridge the approximated solution $\boldsymbol{u}^{t}$ in Algorithm 1 with the optimal solution to the hindsight problem by two LPs, $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$ and $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$. Specifically, let $\mathcal{Y}(\boldsymbol{b},\boldsymbol{d})$ denote the set of optimal solutions to $\phi(\boldsymbol{b},\boldsymbol{d})$ and
$\mathfrak{S}(\boldsymbol{b},\boldsymbol{d},j)=\max_{\boldsymbol{y}\in\mathcal{Y}(\boldsymbol{b},\boldsymbol{d})}y_{j},$
which selects the largest value of $y_{j}$ among all optimal solutions to $\phi(\boldsymbol{b},\boldsymbol{d})$. Note that $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ implies that $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},1)=0$ by Proposition 1. We now show that $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$ under the good event.
Proposition 4.
Given the AIR policy with the time set $\mathcal{T}$ with $\alpha\in(0,1)$ and $\beta\in(\frac{1}{2},1)$, there exist two constants $c_{5}$ and $c_{6}$ independent of $T$ such that when $t\in[c_{5},T-c_{6}]$, under the good event, we have
  1. (i)
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=0$.
  2. (ii)
    $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$.
According to Proposition 4, for most periods, under the good event, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j^{t}}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=0$, resulting in $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$. Therefore, the probability $\mathbb{P}\left(\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})>0\right)$ is upper bounded by the bad event probability. Finally, with the above four steps, we are prepared to prove Theorem 1 by bounding the right-hand side of (5).
$\text{Reg}^{\mathcal{A}}(T)$ $\leq\sum_{t=c_{5}}^{T-c_{6}}r_{\phi}\mathbb{P}\left(\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})>0\right)+(c_{5}+c_{6})r_{\phi}$
$\leq\sum_{t=c_{5}}^{T-c_{6}}r_{\phi}\left(\frac{c_{3}}{(T-t+1)^{2}}+\frac{c_{4}}{(t-1)^{2}}\right)+(c_{5}+c_{6})r_{\phi}$
$\leq\left(\frac{\pi^{2}}{6}(c_{3}+c_{4})+c_{5}+c_{6}\right)r_{\phi},$
which is independent of $T$. Thus, Theorem 1 is proved.
2.4 Finite Resolving
In the above analysis, we establish the constant regret bound when the number of resolvings is $\mathcal{O}(\log\log T)$. Although the number $\mathcal{O}(\log\log T)$ is nearly a constant, it increases in the time horizon, which may still prevent its application to large-scale or time-sensitive problems. A natural question that follows is, what if we are only allowed to solve LPs a finite number of times? In this subsection, we unveil a more detailed relation between resolving frequency and algorithm performance by considering the case when the number of resolvings is a finite number $M$.
In this case, we need to adjust the resolving times for learning and approximation. Given the number $M\geq 2$, we define the finite-resolving schedule $\mathcal{T}^{F}(M)=\mathcal{T}^{F}_{L}(M)\cup\mathcal{T}^{F}_{A}(M)$ as
$\mathcal{T}^{F}_{L}(M)=\left\{\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil,\left\lceil\frac{T}{2}\right\rceil\right\}$
and
$\mathcal{T}^{F}_{A}(M)=\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{M-2}}\right\rceil\right\}.$
To facilitate understanding, we illustrate the resolving schedule in Figure 2. As Figure 2 shows, when the number of resolvings is restricted, we should invest more computational power to the approximation set, and the learning set only needs two time points. More specifically, we will not solve the fluid model until we get enough samples to derive a relatively accurate estimation of $\boldsymbol{p}$, i.e., at the first time point $\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil$. Then, at time point $\lceil T/2\rceil$, we solve the fluid model with a more accurate estimation based on the collected samples. As for the approximation set, in order to correct the approximation error in time, the smaller time points (further away from $T$) are more crucial. Therefore, we keep smaller time points when the number of resolvings is limited.
[…]
Figure 2: Illustration of Finite-Resolving Time Set $\mathcal{T}^{F}(M)=\mathcal{T}^{F}_{L}(M)\cup\mathcal{T}^{F}_{A}(M)$.
We have the following theorem regarding the performance of the algorithm.
Theorem 2.
For $M\geq 2$, given the finite-resolving schedule $\mathcal{T}^{F}(M)$ with $\beta\in(\frac{1}{2},1)$ the regret of the AIR policy is $\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-2}})$.
The proof of Theorem 2 is similar to Theorem 1 and is given in Appendix B.7. As we prove in Section 2.3, for any period $t$, if there exists resolving times $T_{k-1}$ and $T_{k}$ (again we use $T_{k}$ to denote the $k$-th resolving time in $\mathcal{T}^{F}(M)$) such that $t\in[T_{k-1},T_{k})$, then we can prove that the revenue loss at period $t$ is $\mathcal{O}\left(\frac{1}{(t-1)^{2}}+\frac{1}{(T-t+1)^{2}}\right)$. Similarly, we can find that each period $t\in[T_{M},\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-2}}\rceil)$ has the same property, resulting in $\mathcal{O}(1)$ regret during $\left[\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil,\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right)$. Then, due to the revenue loss during the remaining periods, the regret is $\mathcal{O}\left(T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right)$.
According to Theorem 2, by setting $\beta=1/2+\epsilon$, the AIR policy can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret by solving LPs $M$ times. For example, we can achieve an $\mathcal{O}(T^{(1/2+\epsilon)^{2}})\approx\mathcal{O}(T^{1/4})$ regret bound by solving LPs only $M=3$ times, which beats the existing regret bound $\mathcal{O}(\sqrt{T})$ for LP-free algorithms without the non-degeneracy assumption. Therefore, Theorem 2 indicates that the AIR policy can achieve outstanding performance with very limited resolvings.
3 Known Arrival Probabilities
In this section, we consider a variation of the problem in which the arrival probabilities $\boldsymbol{p}$ are known at the beginning. Such cases are widely studied in the network revenue management (NRM) literature. We show that the AIR policy can be easily modified to adapt to this case and achieve a better performance guarantee due to the distribution information.
In the following, we first propose a modified algorithm, which we call the Argmax with Infrequent Resolving and Known Probabilities (AIR-KP) algorithm.
Algorithm 2 Argmax with Infrequent Resolving and Known Probabilities (AIR-KP) Policy
Input: Time set $\mathcal{T}^{\mathcal{K}}=\{T_{1},T_{2},T_{3},\dots,T_{|\mathcal{T}^{\mathcal{K}}|}\}$.
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$, $\boldsymbol{u}^{1}\leftarrow\boldsymbol{0}$ and $\boldsymbol{d}^{1}\leftarrow\boldsymbol{0}$.
for $t=1,2,3,\dots,T$ do
     if $t\in\mathcal{T}$ then $\triangleright$ Infrequent resolving
         Solve the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\boldsymbol{p})$ and obtain its optimal solution $\boldsymbol{y}^{*}$.
         Set $u^{t}_{j}\leftarrow y^{*}_{j}$ and $d^{t}_{j}\leftarrow p_{j}(T-t+1)$ for any $j$.
     end if
     Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
         Accept the request.
         Set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$. $\triangleright$ Update the remaining capacity
         Set $\boldsymbol{u}^{t+1}\leftarrow\boldsymbol{u}^{t}-\boldsymbol{e}_{j}$. $\triangleright$ Approximate the optimal solution
     else
         Reject the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}$.
     end if
     Set $\boldsymbol{d}^{t+1}\leftarrow\boldsymbol{d}^{t}-\boldsymbol{e}_{j}$$\triangleright$ Approximate the future demand
end for
Algorithm 2 is same as Algorithm 1 except that we replace the empirical estimation $\hat{\boldsymbol{p}}^{t}$ with the known arrival probabilities $\boldsymbol{p}$. Next, we present the resolving schedule.
Resolving Schedule. Given the arrival probabilities at the beginning, we can drop the learning time set $\mathcal{T}_{L}$ in $\mathcal{T}$ and add an initial solving at period 1. That is, the resolving schedule for the known-probability case is
$\mathcal{T}^{\mathcal{K}}=\{1\}\cup\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{K_{A}}}\right\rceil\right\},$
where $K_{A}=\lceil\log_{\frac{1}{\beta}}\log_{3}T\rceil$ and $\beta\in(\frac{1}{2},1)$. For ease of understanding, we illustrate the schedule in Figure 3.
[…]
Figure 3: Illustration of Resolving Schedule $\mathcal{T}^{\mathcal{K}}$ for Known-Probability Case.
We show the regret bound of the AIR-KP policy in the following theorem.
Theorem 3 (Regret Bound for NRM).
Given the arrival probabilities and the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(\frac{1}{2},1)$, the regret of the AIR-KP policy is $\mathcal{O}(1)$.
The idea of the proof of Theorem 3 is similar to that of Theorem 1, which is described in Section 2.3. We present the detailed proof in Appendix B.8. According to Theorem 3, for the known-probability case, the AIR-KP policy can guarantee a constant regret bound for the NRM problem under the resolving schedule $\mathcal{T}^{\mathcal{K}}$. Note that the set $\mathcal{T}^{\mathcal{K}}$ is smaller than $\mathcal{T}$ because the learning time set is dropped, but the order is still $O(\log\log T)$. In other words, the distribution information can reduce the resolving times, but the order regarding the time horizon $T$ stays the same.
3.1 Finite Resolving
Similar to the base model, in the following, we consider the known-probability case with a finite number of resolvings. Similar to the case with unknown probabilities, we keep the smaller time points in the approximation set, that is
$\mathcal{T}^{\mathcal{K},F}(M)=\{1\}\cup\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{M-1}}\right\rceil\right\}.$
In the following theorem, we provide a better regret bound compared with Section 2.4.
Theorem 4.
Given the resolving time set $\mathcal{T}^{\mathcal{K},F}(M)$ with $\beta\in(\frac{1}{2},1)$, the regret of the AIR-KP policy is $\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-1}})$.
According to Theorem 4, by setting $\beta=1/2+\epsilon$, we can guarantee an $\mathcal{O}(T^{(1/2+\epsilon)^{M}})\approx\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret by solving LPs $M$ times. The result is close to Theorem 6.1 in Jasin and Kumar (2012), which shows that $M$ LP solving times can induce an $\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret bound under the non-degeneracy assumption. However, different from Jasin and Kumar (2012), our resolving schedule also works for the degenerate case.
3.2 Discussion on Resolving Schedules
In this subsection, in order to show the power of our proof framework, we revisit several resolving schedules proposed in the literature for the known-probability case, i.e., Jasin and Kumar (2012) and Bumpensanti and Wang (2020), and provide corresponding modified schedules for the unknown-probability case. With similar proofs, we can show that the AIR-KP policy is no worse than the proposed policies in Jasin and Kumar (2012) and Bumpensanti and Wang (2020).
  1. 1.
    Periodic Resolving in Jasin and Kumar (2012). In Jasin and Kumar (2012), the authors propose a resolving algorithm for the known-probability case with resolving schedule being
    $\mathcal{T}^{\mathcal{K},P}(\omega)=\left\{1,1+\omega,1+2\omega,\dots,1+K_{P}\omega\right\},$
    with $K_{P}=\left\lfloor\frac{T-1}{\omega}\right\rfloor$. In this case, for $t\in[T_{k-1},T_{k})$, we have $T_{k-1}\geq t-\omega$. Given the resolving schedule $\mathcal{T}^{\mathcal{K},P}(\omega)$, the AIR-KP policy guarantees a constant regret bound for the known-probability case.
    Lemma 3.
    Given the arrival probabilities and the resolving time set $\mathcal{T}^{\mathcal{K},P}(\omega)$, the regret of the AIR-KP policy is $\mathcal{O}(\sqrt{\omega}\log\omega)=\tilde{\mathcal{O}}(\sqrt{\omega})$, which is independent of $T$.
    According to Lemma 3, as the resolving interval $\omega$ increases, the regret bound increases in an $\tilde{\mathcal{O}}(\sqrt{\omega})$ order, which is close to the $\mathcal{O}(\sqrt{\omega})$ regret bound in Jasin and Kumar (2012). Moreover, in the following proposition, the periodic resolving schedule can be directly applied to the unknown-probability case.
    Proposition 5.
    Given the resolving time set $\mathcal{T}^{\mathcal{K},P}(\omega)$, the regret of the AIR policy is $\mathcal{O}(\omega)$, which is independent of $T$.
    The result in Proposition 5 with $\omega=1$ directly improves the regret in the degenerate case from $\tilde{\mathcal{O}}(\sqrt{T})$ in Chen et al. (2024) to $\mathcal{O}(1)$.
  2. 2.
    Midpoint Resolving in Jasin and Kumar (2012). In Jasin and Kumar (2012), the authors also consider the following resolving schedule:
    $\mathcal{T}^{\mathcal{K},M}=\{1\}\cup\left\{\left\lceil T-T/2\right\rceil,\left\lceil T-T/2^{2}\right\rceil,\dots,\left\lceil T-T/2^{K_{M}}\right\rceil\right\}$
    with $K_{M}=\lceil\log_{2}T\rceil$. In this case, for $t\in[T_{k-1},T_{k})$, we have $T-T_{k-1}\leq 2(T-t+1)$. Given the resolving schedule $\mathcal{T}^{\mathcal{K},M}$, our AIR policy guarantees the following regret bound.
    Lemma 4.
    Given the arrival probabilities and the resolving time set $\mathcal{T}^{\mathcal{K},M}$, the regret of the AIR-KP policy is $\mathcal{O}(1)$.
    According to Lemma 4, given the resolving schedule $\mathcal{T}^{\mathcal{K},M}$, the AIR-KP policy can guarantee a constant bound for the known-probability case. However, for the unknown-probability case, since the empirical estimation at the beginning is inaccurate and the second resolving time is $\lceil\frac{T}{2}\rceil$, the regret of the AIR policy will be $\mathcal{O}(T)$. In order to extend such a schedule to fit the unknown-probability case, we will supplement the midpoint resolving schedule with a learning set similar to $\mathcal{T}_{L}$ in the base model. Specifically, we construct a midpoint resolving time set
    $\mathcal{T}^{M}=\left\{\left\lceil T/2^{2}\right\rceil,\left\lceil T/2^{3}\right\rceil,\dots,\left\lceil T/2^{K_{M}}\right\rceil\right\}\bigcup\mathcal{T}^{\mathcal{K},M}.$
    In this case, we have $T-T_{k-1}\leq 2(T-t+1)$ and $T_{k-1}\geq\frac{t-1}{2}$. Under this resolving schedule, our AIR policy can guarantee a constant regret bound.
    Proposition 6.
    Given the resolving time set $\mathcal{T}^{M}$, the regret of the AIR policy is $\mathcal{O}(1)$.
    Note that Jasin (2015) proposes a midpoint-resolving algorithm which can achieve an $\mathcal{O}(\log^{2}T)$ regret under the non-degeneracy assumption. Our result in Proposition 6 provide a way to improve the regret bound even without the non-degeneracy assumption.
  3. 3.
    Infrequent Resolving in Bumpensanti and Wang (2020). In Bumpensanti and Wang (2020), the authors propose a resolving algorithm for the known-probability case with the resolving schedule being the same as $\mathcal{T}^{\mathcal{K}}$ with $\beta=5/6$. According to Theorem 3, for $\beta\in(\frac{1}{2},1)$, our AIR-KP policy can always guarantee a constant regret bound. Moreover, in order to modify such schedules to fit the unknown-probability case, we need to add a learning time set $\mathcal{T}_{L}$ specified in Section 2.2.
Before we close this section, we discuss some extensions regarding the arrival process. First, we consider the case in which the arrival process is a non-stationary process with arrival probabilities $\{p_{j}^{t}:j\in[n],t\in[T]\}$. Similar to Zhu et al. (2023), given the scaling factor $\gamma$, we state the arrival probabilities in the asymptotic regime as $p_{j}^{t}(\gamma)=p_{j}^{\lceil t/\gamma\rceil}$ and $T(\gamma)=\gamma T$. If $\max\{t:p_{j}^{t}>0\}$ is different for different $j$’s, then Zhu et al. (2023) show that the regret is lower bounded by $\Omega(\sqrt{\gamma})$. Otherwise, using similar techniques as in our analysis, we still have a constant bound for the regret.
Second, we consider the case in which the arrival probabilities are modulated by an irreducible, aperiodic and finite-state Markov process with a stationary distribution $\boldsymbol{\pi}$. In this case, we can use $\sum_{\boldsymbol{p}}\pi_{\boldsymbol{p}}p_{j}$ as the arrival probability of type-$j$ customers. Then, due to the exponential convergence of Markov chains (see, e.g., Theorem 15.0.1 in Meyn and Tweedie 2012), the regret can be proved to be $\mathcal{O}(1)$ with a similar proof.
4 Numerical Experiments
In this section, we conduct numerical experiments to compare the performance of different policies. Since the exact problem (2) is intractable, we replace $V^{*}(T)$ in (3) by $\mathbb{E}\left[\phi(\boldsymbol{b}^{1},\boldsymbol{Z}^{1})\right]$ as a benchmark to evaluate policy performance in the following numerical experiments, consistent with the theoretical analysis of regret.
4.1 OLP Policy Comparison
In the following, we numerically compare the AIR policy with several OLP algorithms as follows (see Appendix C.1 for the detailed descriptions of these algorithms):
  1. 1.
    AIR Policy. We set the parameters in the resolving schedule as $\alpha=\beta=0.7$.
  2. 2.
    Argmax with Frequent Resolving (AFR) Algorithm. We study the case when the LP is resolved per period, which is similar to Algorithm 3 in Xie et al. (2023).
  3. 3.
    Adaptive Allocation (ADA) Policy. We implement Algorithm 1 in Chen et al. (2024).
  4. 4.
    Simple and Fast (SFA) Policy. We implement Algorithm 5 in Li et al. (2020) with stepsize $\gamma_{t}=1/\sqrt{t}$.
  5. 5.
    Decoupling Learning and Decision (DLD) Policy. We implement Algorithm 2 in Gao et al. (2024) (the parameters are $T_{e}=\lfloor T^{2/3}\rfloor$, $\alpha_{e}=T^{-1/3}$ and $\alpha_{p}=T^{-2/3}$) with $\mathcal{A}_{L}$ being the subgradient algorithm with stepsize $\gamma_{L,t}=1/t$ and $\mathcal{A}_{D}$ being the subgradient algorithm with stepsize $\alpha_{e}$ before time $T_{e}$ and $\alpha_{p}$ after that. In order to satisfy the budget constraints, the modified algorithm will always reject a request if the request cannot be fulfilled, i.e., $\boldsymbol{A}_{j^{t}}>\boldsymbol{b}^{t}$.
  6. 6.
    Budget-Updating Fast (BUF) Policy. We implement Algorithm 5 in Ma et al. (2024).
In the following, we compare the performance of the above algorithms. For each parameter set, we run 2,000 simulations. To gain deeper insights, we first focus on the single resource problem similar to the one studied in Bumpensanti and Wang (2020). Specifically, we consider two types of customers and one type of resource. The arrival probabilities of either customer type is $0.5$. The rewards of the two types of customers are $r_{1}=2$ and $r_{2}=1$, and each customer consumes one unit of resource. The budget per period is $\rho$.
First, in Figure 4, we fix the time horizon $T$ as $50,000$, and test the algorithms as the budget factor $\rho$ changes. According to Figure 4, the regret of most policies is large when $\rho$ is close to 0.5. Due to the setup of the single-resource problem, the fluid problem at time $1$ is degenerate when $\rho=0.5$. Therefore, the degenerate case is the hardest case for most policies. However, both the AIR policy and the AFR policy guarantee a constant regret in this case.
[…]
Figure 4: Regret under Different Policies as Functions of $\rho$ when $m=1$, $n=2$, $r_{1}=2$, $r_{2}=1$, $p_{1}=p_{2}=0.5$, $T=50,000$ and $\alpha=\beta=0.7$.
We then focus on the degenerate case and numerically test the above algorithms as the time horizon $T$ increases. Specifically, given $m=10$ and $n=2$, we randomly generate the parameters $\boldsymbol{A}$, $\boldsymbol{p}$ and $\boldsymbol{r}$, and choose $\boldsymbol{\rho}$ such that the optimal solution is degenerate (see Appendix C.2 for detailed parameters). We run 200 simulations for each set of parameters and summarize the regret and the computation time of the above policies in Table 3. In addition, we present the detailed resolving schedule in Table 4.
Time Horizon Regret Computation Time (s) # of Resolvings
AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA SFA* DLD* BUF* AIR
2500 2.5 1.5 7.7 45.6 62.3 48.3 0.062 13.3 13.3 0.002 0.003 0.002 13
5000 2.2 1.2 10.5 57.6 82.6 59.0 0.064 26.4 26.6 0.004 0.005 0.004 13
7500 2.2 1.6 12.0 66.6 96.4 65.7 0.065 39.4 39.3 0.005 0.007 0.005 13
10,000 2.2 1.4 13.2 74.4 109.7 72.5 0.066 54.4 54.2 0.007 0.009 0.007 13
12,500 2.1 1.2 14.3 80.9 118.8 76.0 0.077 66.6 66.5 0.009 0.012 0.009 15
15,000 2.2 1.3 15.3 86.8 128.1 79.7 0.086 73.3 73.2 0.011 0.015 0.011 15
17,500 2.2 1.1 16.6 92.1 136.0 82.9 0.083 84.4 84.3 0.012 0.016 0.012 15
20,000 2.1 1.0 17.4 97.0 141.6 85.9 0.084 103.2 103.1 0.014 0.019 0.013 15
100,000 2.2 - [^4] The case is not studied due to the extremely long computation time. - 192.0 260.1 126.6 0.141 - - 0.080 0.105 0.078 15
200,000 2.1 - - 260.2 330.9 151.6 0.204 - - 0.169 0.219 0.163 15
300,000 2.1 - - 313.8 379.2 166.1 0.231 - - 0.220 0.291 0.213 15
Table 3: Regret and Computation Time of Different Policies when $m=10$, $n=2$ and $\alpha=\beta=0.7$. Superscript * Marks LP-free Policies.
From Table 3, we observe several interesting phenomena regarding the regret. First, the regret of either LP-free policy (SFA, DLD or BUF) increases in the time horizon. In contrast, the regret of both the AFR policy and the AIR policy remains constant, highlighting the effectiveness of LP resolving. Second, the regret of the ADA policy, which also solves LPs per period, also increases in the time horizon. This implies that the choice of the base policy or the interpretation of the fluid model’s solution is important. Lastly, the computational time of the AIR policy is almost neglectable, while the performance of the AIR policy is close to the AFR policy which solves LP in every period, and is much better than those LP-free policies. For example, when $T=20,000$, the AIR policy, which solves LPs only 15 times (see Table 4 for the detailed resolving schedule), achieves a regret in the same order as the AFR policy with only about 0.1% of the computational time. This suggests that an appropriate resolving schedule is important.
Time Horizon $\mathcal{T}_{L}$ $\mathcal{T}_{A}$
2500 3 4 7 15 47 240 1250 2261 2454 2486 2494 2497 2498
5000 3 5 8 19 65 389 2500 4621 4936 4982 4993 4996 4998
7500 3 5 9 22 80 516 3750 6985 7421 7479 7492 7496 7498
10,000 3 5 10 24 92 631 5000 9370 9909 9977 9991 9996 9998
12,500 3 4 5 10 26 102 738 6250 11763 12399 12475 12491 12496 12497 12498
15,000 3 4 6 11 28 112 839 7500 14162 14889 14973 14990 14995 14997 14998
17,500 3 4 6 11 29 120 934 8750 16567 17381 17472 17490 17495 17497 17498
20,000 3 4 6 11 30 129 1025 10000 18976 19872 19971 19990 19995 19997 19998
100,000 3 4 7 16 52 282 3163 50000 96838 99719 99949 99985 99994 99997 99998
200,000 3 5 8 19 66 396 5138 100000 194863 199605 199935 199982 199993 199996 199998
300,000 3 5 9 21 76 483 6824 150000 293177 299518 299925 299980 299992 299996 299998
Table 4: Resolving Schedule $\mathcal{T}$ with $\alpha=\beta=0.7$.
Lastly, in order to identify the impact of $n$, we consider the case when $m=10$ and $n=50$, and compare the AIR policy with LP-free policies. For each parameter set, we run 200 simulations. The comparison results are illustrated in Figure 5. As Figure 5 shows, the regret of the AIR policy is still a small constant. Note that in this case, we set $\alpha=\beta=0.9$ and the number of resolvings is no greater than 43 for $T\leq 30,000$. The reason is that as the number of customer types increases, the optimal solution of the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}})$ changes more significantly and frequently and hence the resolving frequency should be slightly increased.
[…]
Figure 5: Regret under Different Policies as Functions of $T$ when $m=10$, $n=50$ and $\alpha=\beta=0.9$.
4.2 Finite Resolving
In this case, we study the AIR policy with the resolving schedule $\mathcal{T}^{F}(3)$ with $\beta=0.7$, referred to as “AIR-3”, and use the degenerate case with $m=10$ and $n=2$ in Section 4.1. For each parameter set, we run 2,000 simulations. In Figure 6, we compare performance of the AIR-3 policy with LP-free policies (e.g., SFA, DLD, BUF). According to Figure 6, the AIR-3 policy can guarantee a low regret by solving LP only three times, illustrating the power of LP resolving.
[…]
Figure 6: Regret under Different Policies as Functions of $T$ when $m=10$ and $n=2$.
4.3 Known-Probability Case
In this subsection, we consider the single-resource case in Section 4.1. We compare three policies: R-PAC policy with per-period resolving in Jasin and Kumar (2012), IRT policy in Bumpensanti and Wang (2020), and our AIR policy with $\beta=5/6$. In this case, the resolving schedule of our policy is the same as that of the IRT policy. In Figure 7, we illustrate the regret $\text{Reg}^{\mathcal{A}}(T)$ as the budget per period $\rho$ changes.
[…]
Figure 7: Regret under Different Policies as Functions of $\rho$ when $T=50,000$ and $\beta=5/6$.
According to Figure 7, the regret bound of the R-PAC policy is large when the initial LP is degenerate ($\rho=0.5$), which agrees with the lower bound $\Omega(\sqrt{T})$ in Bumpensanti and Wang (2020). In contrast, both the IRT policy and our R-AIR policy can deal with the degeneracy well and hence keep the constant bound. Note that both the IRT and the AIR policies only solves LPs 14 times (with the same schedule) while the R-PAC policy solves LPs 50,000 times, which highlights the effectiveness of the resolving schedule.
5 Concluding Remarks
In this paper, we investigate the OLP problem under the random input assumption, assuming a finite-support underlying distribution. We propose an infrequent resolving algorithm that guarantees constant regret by solving LPs $\mathcal{O}(\log\log T)$ times. This algorithm strikes a superior balance between algorithm performance (i.e., regret) and computation efficiency (i.e., resolving frequency). Compared to LP-based algorithms, we significantly reduce the number of resolvings without sacrificing algorithm performance too much; compared to LP-free algorithms, we substantially improve performance by a slight increase in resolving frequency. Moreover, given that the number of resolvings is a finite number $M$, we design a resolving schedule such that our policy guarantees an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret bound. Furthermore, when the arrival probabilities are known at the beginning, we can also guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times.
From a methodological point of view, our work introduces a novel framework for proving regret bounds of infrequent resolving algorithms. The design of the resolving schedule highlights the importance of resolving at both the beginning and the end of the time horizon. These insights may be helpful in designing infrequent resolving algorithms for other problems.
Appendix
Appendix A Concentration Inequality
In this paper, we mainly use the Hoeffding’s inequality in Hoeffding (1963). To be self-contained, we present the inequality in the following lemma.
Lemma 5 (Hoeffding’s Inequality).
Let $V_{1},V_{2},\dots,V_{k}$ be i.i.d. Bernoulli random variables with mean $\bar{v}$. Then, we have
$\mathbb{P}\left(\sum_{\ell=1}^{k}V_{\ell}-k\bar{v}\geq\sqrt{k\log\xi}\right)\leq\frac{1}{\xi^{2}}$
$\mathbb{P}\left(\sum_{\ell=1}^{k}V_{\ell}-k\bar{v}\leq-\sqrt{k\log\xi}\right)\leq\frac{1}{\xi^{2}}$
$\mathbb{P}\left(\left|\sum_{\ell=1}^{k}V_{\ell}-k\bar{v}\right|\geq\sqrt{k\log\xi}\right)\leq\frac{2}{\xi^{2}}.$
Appendix B Omitted Proofs
In this section, we provide omitted proofs in the maintext.
B.1 Proof of Lemma 2
Let $\mu^{*}$ denote the optimal policy of $V^{*}(T)$. First, for any sample path $\omega$ with the demand $\boldsymbol{Z}^{1}(\omega)$, let $\bar{y}_{j}(\omega)=\sum_{t=1}^{T}x_{t,j}^{\mu^{*}}(\omega)$ for each $j$. Then, it can be verified that $\bar{\boldsymbol{y}}(\omega)$ is always a feasible solution to the problem $\phi(\boldsymbol{b}^{1},\boldsymbol{Z}^{1}(\omega))$ because of the feasibility constraints in (2). Then, we have $\mathbb{E}\left[\mathbb{E}[\phi(\boldsymbol{b}^{1},\boldsymbol{Z}^{1}):\omega]\right]\geq\mathbb{E}[\boldsymbol{r}^{\mathsf{T}}\bar{\boldsymbol{y}}(\omega)]=V^{*}(T)$. ∎
B.2 Proof of Proposition 1
In this proof, we fix $(\boldsymbol{b},\boldsymbol{Z},j)$ with $\boldsymbol{b}\geq\boldsymbol{A}_{j}x$ and $\boldsymbol{Z}\geq\boldsymbol{e}_{j}$, and define $\bar{a}:=\max_{i,j}a_{ij}$ and $\bar{r}=\max_{j}r_{j}$. First, according to Theorem 2.4 in Mangasarian and Shiau (1987), for any optimal solution $\boldsymbol{y}_{1}^{*}$ of $\phi(\boldsymbol{b},\boldsymbol{Z})$, there exists an optimal solution $\boldsymbol{y}_{2}^{*}$ of $\phi(\boldsymbol{b}-\boldsymbol{A}_{j},\boldsymbol{Z}-\boldsymbol{e}_{j})$ such that
$\|\boldsymbol{y}_{1}^{*}-\boldsymbol{y}_{2}^{*}\|_{\infty}\leq\kappa_{1}\cdot\max\{\max_{i}a_{ij},1\}\leq\kappa_{1}\cdot\max\{\bar{a},1\},$
where $\kappa_{1}$ only depends on the matrix $\boldsymbol{A}$. Then, we have
$\Delta(\boldsymbol{b},\boldsymbol{Z},j,x)=\phi(\boldsymbol{b},\boldsymbol{Z})-\phi(\boldsymbol{b}-x\boldsymbol{A}_{j},\boldsymbol{Z}-\boldsymbol{e}_{j})-xr_{j}\leq\sum_{j=1}^{n}\kappa_{1}r_{j}\cdot\max\{\bar{a},1\}\leq n\kappa_{1}\bar{r}\cdot\max\{\bar{a},1\}.$
Therefore, we have $\Delta(\boldsymbol{b},\boldsymbol{Z},j,x)\leq r_{\phi}:=n\kappa_{1}\bar{r}\cdot\max\{\bar{a},1\}$, which is independent of $T$.
Let $\boldsymbol{y}^{*}$ be an optimal solution of $\phi(\boldsymbol{b},\boldsymbol{Z})$ with $y_{j}^{*}\geq 1$. Then, we prove that $\boldsymbol{y}^{*}-\boldsymbol{e}_{j}$ is an optimal solution to $\phi(\boldsymbol{b}-\boldsymbol{A}_{j},\boldsymbol{Z}-\boldsymbol{e}_{j})$. First, it is obvious that $\boldsymbol{y}^{*}-\boldsymbol{e}_{j}$ is a feasible solution. Second, suppose there exists a feasible solution $\tilde{\boldsymbol{y}}$ of $\phi(\boldsymbol{b}-\boldsymbol{A}_{j},\boldsymbol{Z}-\boldsymbol{e}_{j})$ such that $\boldsymbol{r}^{\mathsf{T}}\tilde{\boldsymbol{y}}>\boldsymbol{r}^{\mathsf{T}}(\boldsymbol{y}^{*}-\boldsymbol{e}_{j})$. Since $\tilde{\boldsymbol{y}}+\boldsymbol{e}_{j}$ is a feasible solution to $\phi(\boldsymbol{b},\boldsymbol{Z})$, we have $\boldsymbol{r}^{\mathsf{T}}(\tilde{\boldsymbol{y}}+\boldsymbol{e}_{j})>\boldsymbol{r}^{\mathsf{T}}\boldsymbol{y}^{*}$, which contradicts with the optimality of $\boldsymbol{y}^{*}$. Therefore, we can deduce that $\boldsymbol{y}^{*}-\boldsymbol{e}_{j}$ is an optimal solution to $\phi(\boldsymbol{b}-\boldsymbol{A}_{j},\boldsymbol{Z}-\boldsymbol{e}_{j})$, implying that $\Delta(\boldsymbol{b},\boldsymbol{Z},j,1)=0$. Similarly, we can prove that $\Delta(\boldsymbol{b},\boldsymbol{Z},j,0)=0$ if $Z_{j}-y_{j}^{*}\geq 1$. ∎
B.3 Proof of Proposition 2
For $t\in[T_{k-1},T_{k})$, we have
$|d_{j}^{t}-Z_{j}^{t}|$ $=\left|\frac{(\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell})\cdot(T-T_{k-1}+1)}{T_{k-1}-1}-\sum_{\ell=T_{k-1}}^{t-1}Y_{j}^{\ell}-\sum_{\ell=t}^{T}Y_{j}^{\ell}\right|$
$=\left|\frac{(\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell})\cdot(T-T_{k-1}+1)}{T_{k-1}-1}-\sum_{\ell=T_{k-1}}^{T}Y_{j}^{\ell}\right|$
$\leq(T-T_{k-1}+1)\left(\left|\frac{(\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell})}{T_{k-1}-1}-p_{j}\right|+\left|p_{j}-\frac{\sum_{\ell=T_{k-1}}^{T}Y_{j}^{\ell}}{T-T_{k-1}+1}\right|\right),$
where the first equality follows from the definition of $d_{j}^{t}$. Note that the arrival process is i.i.d. across time, by Hoeffding’s inequality, it holds that
$\mathbb{P}\left(|d_{j}^{t}-Z_{j}^{t}|\geq(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}+\sqrt{(T-T_{k-1}+1)\log(T-t+1)}\right)$
$\leq$ $\mathbb{P}\left(\left|\frac{(\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell})}{T_{k-1}-1}-p_{j}\right|\geq\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}\right)+\mathbb{P}\left(\left|p_{j}-\frac{\sum_{\ell=T_{k-1}}^{T}Y_{j}^{\ell}}{T-T_{k-1}+1}\right|\geq\sqrt{\frac{\log(T-t+1)}{T-T_{k-1}+1}}\right)$
$\leq$ $\frac{2}{(t-1)^{2}}+\frac{2}{(T-t+1)^{2}}.$
Similarly, we have $|d_{j}^{t}-Z_{j}^{t}|\leq(T-T_{k-1}+1)\sqrt{\frac{\log(T-t+1)}{T_{k-1}-1}}+\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$ with probability at least $1-\frac{4}{(T-t+1)^{2}}$. ∎
B.4 Proof of Lemma 1
Given $t\in[T_{k-1},T_{k})$, there exist $n_{1}$ and $n_{2}$ such that $T_{k-1}\geq\max\{\lceil T^{\alpha^{n_{1}+1}}\rceil,\lceil T-T^{\beta^{n_{2}}}\rceil\}$ and $T_{k}\leq\max\{\lceil T^{\alpha^{n_{1}}}\rceil,\lceil T-T^{\beta^{n_{2}+1}}\rceil\}$. Then, we have
$(t-1)^{\alpha}\leq(T_{k}-1)^{\alpha}\leq(\lceil T^{\alpha^{n_{1}}}\rceil-1)^{\alpha}\leq(T^{\alpha^{n_{1}}})^{\alpha}=T^{\alpha^{n_{1}+1}}\leq T_{k-1}$
$(T-t+1)^{1/\beta}\geq(T-T_{k}+1)^{1/\beta}\geq(T-\lceil T-T^{\beta^{n_{2}+1}}\rceil+1)^{1/\beta}\geq(T^{\beta^{n_{2}+1}})^{1/\beta}=T^{\beta^{n_{2}}}\geq T-T_{k-1}.$
B.5 Proof of Proposition 3
We prove the two statements in Proposition 3 one by one. Before proceeding, we first simplify $d_{j}^{t}-(T-t+1)p_{j}$ as follows:
$d_{j}^{t}-(T-t+1)p_{j}$ $=\frac{(\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell})\cdot(T-T_{k-1}+1)}{T_{k-1}-1}-\sum_{\ell=T_{k-1}}^{t-1}Y_{j}^{\ell}-(T-t+1)p_{j}$
$=(T-T_{k-1}+1)\left(\frac{\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell}}{T_{k-1}-1}-p_{j}\right)+(t-T_{k-1})\left(p_{j}-\frac{\sum_{\ell=T_{k-1}}^{t-1}Y_{j}^{\ell}}{t-T_{k-1}}\right).$
(i) Consider the time period $t\in[3,T-3]$ such that there exist $T_{k-1}$ and $T_{k}$ satisfying $t\in[T_{k-1},T_{k})$. There are two possible cases:
  1. (a)
    When $T_{k-1}<\lceil\frac{T}{2}\rceil$, similar to the proof of Proposition 2, by Hoeffding’s inequality, we have
    $\mathbb{P}\left(d_{j}^{t}-(T-t+1)p_{j}\leq-(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\sqrt{(t-T_{k-1})\log(T-t+1)}\right)$
    $\leq$ $\mathbb{P}\left(\frac{(\sum_{\ell=1}^{T_{k-1}-1}Y_{j}^{\ell})}{T_{k-1}-1}-p_{j}\leq-\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}\right)+\mathbb{P}\left(p_{j}-\frac{\sum_{\ell=T_{k-1}}^{t-1}Y_{j}^{\ell}}{t-T_{k-1}}\leq-\sqrt{\frac{\log(T-t+1)}{t-T_{k-1}}}\right)$
    $\leq$ $\frac{1}{(t-1)^{2}}+\frac{1}{(T-t+1)^{2}}.$
    Since $\lceil\frac{T}{2}\rceil\in\mathcal{T}$ and $T_{k-1}<\lceil\frac{T}{2}\rceil$, we have $t-T_{k-1}\leq T-T_{k-1}+1\leq T\leq 2(T-T_{k}+1)\leq 2(T-t+1)$. Then, according to Lemma 1, with probability at least $1-\frac{1}{(t-1)^{2}}-\frac{1}{(T-t+1)^{2}}$, it holds
    $d_{j}^{t}\geq(T-t+1)p_{j}-2(T-t+1)\sqrt{\frac{\log(t-1)}{(t-1)^{\alpha}-1}}-\sqrt{2(T-t+1)\log(T-t+1)}.$
    Since $\alpha>0$, there exists a constant $\eta_{j}$ such that when $t\geq\eta_{j}$, we have $2\sqrt{\frac{\log(t-1)}{(t-1)^{\alpha}-1}}\leq\frac{p_{j}}{4}$. Moreover, there exists a constant $\eta^{\prime}_{j}$ such that when $t\leq T-\eta^{\prime}_{j}$, we have
    $\frac{p_{j}}{4}(T-t+1)-\sqrt{2(T-t+1)\log(T-t+1)}\geq 0\text{ and }\frac{p_{j}(T-t+1)}{2}\geq 2.$
    Therefore, when $t\in[\eta_{j},\min\{T-\eta^{\prime}_{j},\lceil\frac{T}{2}\rceil\}]$, we have
    $d_{j}^{t}\geq\frac{3p_{j}(T-t+1)}{4}-\sqrt{2(T-t+1)\log(T-t+1)}\geq\frac{p_{j}(T-t+1)}{2}\geq 2$
    with probability at least $1-\frac{1}{(t-1)^{2}}-\frac{1}{(T-t+1)^{2}}$.
  2. (b)
    When $T_{k-1}\geq\lceil\frac{T}{2}\rceil$, according to the Hoeffding’s inequality, we have
    $d_{j}^{t}\geq(T-t+1)p_{j}-(T-T_{k-1}+1)\sqrt{\frac{\log(T-t+1)}{T_{k-1}-1}}-\sqrt{(t-T_{k-1})\log(T-t+1)},$
    with probability at least $1-\frac{2}{(T-t+1)^{2}}$. Since $T_{k-1}\geq\lceil\frac{T}{2}\rceil$, we have $T_{k-1}-1\geq\frac{T-T_{k-1}+1}{2}$. According to Lemma 1, we have
    $d_{j}^{t}$ $\geq(T-t+1)p_{j}-(T-T_{k-1}+1)\sqrt{\frac{2\log(T-t+1)}{T-T_{k-1}+1}}-\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $=(T-t+1)p_{j}-(\sqrt{2}+1)\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq(T-t+1)p_{j}-(\sqrt{2}+2)(T-t+1)^{\frac{1}{2\beta}}\sqrt{\log(T-t+1)}.$
    Since $\frac{1}{2\beta}<1$, there exists a constant $\eta^{\prime\prime}_{j}$ such that when $t\leq T-\eta^{\prime\prime}_{j}$, we have $d_{j}^{t}\geq\frac{p_{j}(T-t+1)}{2}\geq 2$ with probability $1-\frac{2}{(T-t+1)^{2}}$.
Let $c_{1}=\max\{\max_{j}\eta_{j}$, 3} and $c_{2}=\max\{\max_{j}\eta^{\prime}_{j},\max_{j}\eta^{\prime\prime}_{j},3\}$, that are independent of $T$. The above proof implies that, when $t\in[c_{1},T-c_{2}]$, we have $\boldsymbol{d}^{t}\geq\frac{p_{j}(T-t+1)}{2}\geq\boldsymbol{2}$ with probability $1-\frac{n}{(T-t+1)^{2}}-\frac{n}{(\min\{t-1,T-t+1\})^{2}}$.
(ii) According to the proof of (i), when $t\in[c_{1},T-c_{2}]\cap[T_{k-1},T_{k})$, we have $\boldsymbol{d}^{t}\geq\boldsymbol{2}$ with high probability. Then, it suffice to prove that if $\boldsymbol{d}^{t}\geq\boldsymbol{2}$, then $\boldsymbol{u}^{t}$ is an optimal solution to the surrogate LP $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$. Given that $\boldsymbol{d}^{t}\geq\boldsymbol{2}$, we have $\boldsymbol{d}^{\ell}\geq\boldsymbol{2}$ when $\ell\in[T_{k-1},t-1]$ due to the monotonicity of $\boldsymbol{d}^{t}$. When $t=T_{k-1}$, the solution $\boldsymbol{u}^{t}$ is an optimal solution solved from $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$. Then, we prove the statement by induction.
Suppose $\boldsymbol{u}^{\tau}$ is an optimal solution to $\phi(\boldsymbol{b}^{\tau},\boldsymbol{d}^{\tau})$ with $\tau\in[T_{k-1},t)$. Since $\boldsymbol{d}^{\tau}\geq\boldsymbol{2}$, if the AIR algorithm accepts the arriving customer of type $j^{t}$, then we have $u^{\tau}_{j^{t}}\geq\frac{1}{2}d^{\tau}_{j^{t}}\geq 1$. Thus, we have $\boldsymbol{u}^{\tau+1}=\boldsymbol{u}^{\tau}-\boldsymbol{e}_{j^{t}}\geq\boldsymbol{0}$ is a feasible solution of $\phi(\boldsymbol{b}^{\tau+1},\boldsymbol{d}^{\tau+1})$. If there exists a feasible solution $\tilde{\boldsymbol{u}}$ of $\phi(\boldsymbol{b}^{\tau+1},\boldsymbol{d}^{\tau+1})$ such that $\boldsymbol{r}^{\mathsf{T}}\tilde{\boldsymbol{u}}>\boldsymbol{r}^{\mathsf{T}}\boldsymbol{u}^{\tau+1}$, then we have $\tilde{\boldsymbol{u}}+\boldsymbol{e}_{j^{t}}$ feasible to $\phi(\boldsymbol{b}^{\tau},\boldsymbol{d}^{\tau})$ and $\boldsymbol{r}^{\mathsf{T}}(\tilde{\boldsymbol{u}}+\boldsymbol{e}_{j^{t}})>\boldsymbol{r}^{\mathsf{T}}\boldsymbol{u}^{\tau}$, which contradicts the optimality of $\boldsymbol{u}^{\tau}$. Then, we can deduce that $\boldsymbol{u}^{\tau+1}$ is an optimal solution to $\phi(\boldsymbol{b}^{\tau+1},\boldsymbol{d}^{\tau+1})$. If the AIR algorithm rejects the arriving customer, we can similarly prove the optimality of $\boldsymbol{u}^{\tau+1}$. Therefore, we can prove that $\boldsymbol{u}^{t}$ is optimal to $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$. ∎
B.6 Proof of Proposition 4
In the following, we always consider the good event. At the beginning, we first bound the difference $|\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})-\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})|$. According to Theorem 2.4 in Mangasarian and Shiau (1987), for any optimal solution $\boldsymbol{y}_{1}$ to $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$, there exists an optimal solution $\boldsymbol{y}_{2}$ to $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$ such that $\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{\infty}\leq\kappa_{2}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}$, where $\kappa_{2}$ is a constant independent of $T$. Therefore, we have
$|\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})-\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})|=|\boldsymbol{r}^{\mathsf{T}}(\boldsymbol{y}_{1}-\boldsymbol{y}_{2})|\leq\left(\sum_{j=1}^{n}r_{j}\right)\|\boldsymbol{y}_{1}-\boldsymbol{y}_{2}\|_{\infty}\leq\kappa_{2}\left(\sum_{j}r_{j}\right)\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}.$
Note that $\mathfrak{S}(\boldsymbol{b},\boldsymbol{d},j)$ can be formulated as an LP:
$\max_{\boldsymbol{y}\geq\boldsymbol{0}}\quad$ $y_{j}$
s.t. $\boldsymbol{r}^{\mathsf{T}}\boldsymbol{y}\geq\phi(\boldsymbol{b},\boldsymbol{d}),$
$\boldsymbol{A}\boldsymbol{y}\leq\boldsymbol{b},$
$\boldsymbol{y}\leq\boldsymbol{d}.$
According to Theorem 2.4 in Mangasarian and Shiau (1987), we have
$\|\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{d}^{t},j)\|_{\infty}\leq\kappa_{3}(\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}+|\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})-\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})|)\leq\kappa_{4}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty},$
where $\kappa_{4}=\kappa_{3}+\kappa_{2}\sum_{j}r_{j}$.
Consider the case when a type-$j$ customer arrives at time $t$ and the AIR policy accepts this request. We then have
$\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$ $\geq\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{d}^{t},j)-\kappa_{4}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}\geq u_{j}^{t}-\kappa_{4}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}$
$\geq\frac{d_{j}^{t}}{2}-\kappa_{4}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{4}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty},$
where the second inequality holds because $\boldsymbol{u}^{t}$ is an optimal solution to $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$, the third inequality is due to the argmax operation in Algorithm 1, and the last inequality is due to Proposition 3.
Given $t\in[T_{k-1},T_{k})$, we discuss two possible cases:
  1. (a)
    When $T_{k-1}<\lceil\frac{T}{2}\rceil$, since $\lceil\frac{T}{2}\rceil\in\mathcal{T}$ and $T_{k-1}<\lceil\frac{T}{2}\rceil$, we have $t\leq\lceil\frac{T}{2}\rceil$ and $T-T_{k-1}+1\leq T\leq 2(T-t+1)$. According to Proposition 2, we have
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\kappa_{4}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-2\kappa_{4}(T-t+1)\sqrt{\frac{\log(t-1)}{(t-1)^{\alpha}-1}}-\kappa_{4}\sqrt{2(T-t+1)\log(T-t+1)}.$
    where the last inequality is because of Lemma 1. Since $\alpha>0$, there exists a constant $\theta_{j}$ such that when $t\geq\theta_{j}$, we have $2\kappa_{4}\sqrt{\frac{\log(t-1)}{(t-1)^{\alpha}-1}}\leq\frac{p_{j}}{8}$. Moreover, there exists a constant $\theta^{\prime}_{j}$ such that when $t\leq T-\theta^{\prime}_{j}$, we have $\frac{(T-t+1)p_{j}}{8}-\kappa\sqrt{2(T-t+1)\log(T-t+1)}\geq 1.$ Therefore, when $t\in[\theta_{j},\min\{\lceil\frac{T}{2}\rceil,T-\theta^{\prime}_{j}\}]$, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq\frac{(T-t+1)p_{j}}{8}-\kappa_{4}\sqrt{2(T-t+1)\log(T-t+1)}\geq 1$.
  2. (b)
    When $T_{k-1}\geq\lceil\frac{T}{2}\rceil$, we have $T_{k-1}-1\geq\frac{T-T_{k-1}+1}{2}$. According to Proposition 2, we have
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-T_{k-1}+1)\sqrt{\frac{\log(T-t+1)}{T_{k-1}-1}}-\kappa_{4}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-T_{k-1}+1)\sqrt{\frac{2\log(T-t+1)}{T-T_{k-1}+1}}-\kappa_{4}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $=$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(\sqrt{2}+1)\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(2+\sqrt{2})(T-t+1)^{\frac{1}{2\beta}}\sqrt{\log(T-t+1)}.$
    where the last inequality is because of Lemma 1. Similar to the above discussions, since $\beta>\frac{1}{2}$, there exists a constant $\theta^{\prime\prime}_{j}$ such that $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$ when $\lceil\frac{T}{2}\rceil\leq t\leq T-\theta^{\prime\prime}_{j}$.
Let $\kappa_{5}^{A}=\max_{j}\theta_{j}$ and $\kappa_{6}^{A}=\max\{\max_{j}\theta^{\prime}_{j},\max_{j}\theta^{\prime\prime}_{j}\}$, which are independent of $T$. When $t\in[\kappa_{5}^{A},T-\kappa_{6}^{A}]\cap[c_{1},T-c_{2}]$, if the algorithm accepts a type-$j$ customer at time $t$, then we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$. Similarly, for the rejection action, we can derive $\kappa_{5}^{R}$ and $\kappa_{6}^{R}$. By setting $c_{5}=\max\{\kappa_{5}^{A},\kappa_{5}^{R},c_{1}\}$ and $c_{6}=\max\{\kappa^{A}_{6},\kappa^{R}_{6},c_{2}\}$, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=0$. Finally, according to Proposition 1, we have $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$. ∎
B.7 Proof of Theorem 2
We discuss different time intervals.
  1. 1.
    For the time interval $\left[1,\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right)$, Algorithm 1 incurs at most $T^{(1/2+\epsilon)\cdot\beta^{M-2}}\cdot r_{\phi}$ regret.
  2. 2.
    For the time interval $\left[\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil,\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right]$, the proofs of Propositions 2, 3 and 4 still hold, and hence the regret at time $t$ can be bounded by $\left(\frac{c_{3}}{(T-t+1)^{2}}+\frac{c_{4}}{(t-1)^{2}}\right)r_{\phi}$.
  3. 3.
    For the time interval $\left[\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil+1,T\right]$, the regret is no greater than $T^{(1/2+\epsilon)\cdot\beta^{M-2}}r_{\phi}$.
To summarize, given the resolving time set $\mathcal{T}^{F}(M)$, the regret can be bounded as
$\text{Reg}^{\mathcal{A}}(T)$ $\leq T^{(1/2+\epsilon)\cdot\beta^{M-2}}\cdot r_{\phi}+\sum_{t=\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil}^{\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil}\left(\frac{c_{3}}{(T-t+1)^{2}}+\frac{c_{4}}{(t-1)^{2}}\right)r_{\phi}+T^{(1/2+\epsilon)\cdot\beta^{M-2}}r_{\phi}$
$\leq 2T^{(1/2+\epsilon)\cdot\beta^{M-2}}r_{\phi}+\frac{\pi^{2}}{6}(c_{3}+c_{4})r_{\phi}=\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-2}}).$
B.8 Proof of Theorem 3
Similar to Lemma 1, we have $T-T_{k-1}\leq(T-t+1)^{1/\beta}$ for any $t\in[T_{k-1},T_{k})$. In this case, the regret decomposition is the same as (5) but the proof is simpler than the unkown-probability case. In the following, we present the main steps of the proof. First, we bound the approximation error, i.e., the difference between $\boldsymbol{d}^{t}$ and $\boldsymbol{Z}^{t}$.
Lemma 6.
Consider the resolving schedule $\mathcal{T}^{\mathcal{K}}$. Given a time $t\in[T_{k-1},T_{k})$, we have $|d_{j}^{t}-Z_{j}^{t}|\leq\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$ with probability larger than $1-\frac{2}{(T-t+1)^{2}}$.
Second, we prove the relationship between $\boldsymbol{u}^{t}$ and a surrogate LP.
Lemma 7.
Consider the known-probability case. For the AIR-KP policy with the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(1/2,1)$, there exists a constant $c_{7}$ independent of $T$ such that when $t\leq T-c_{7}$, with probability larger than $1-\frac{n}{(T-t+1)^{2}}$, we have
  1. 1.
    $d_{j}^{t}\geq\frac{p_{j}(T-t+1)}{2}\geq 2$ for any $j$.
  2. 2.
    $\boldsymbol{u}^{t}$ is an optimal solution to the surrogate LP $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$.
Third, we bound the difference between $\boldsymbol{u}^{t}$ and the hindsight solution, and then prove the regret bound. The good event is defined as the case when conditions in both Lemmas 6 and 7 are satisfied, and the probability is no less than $1-\frac{c_{8}}{(T-t+1)^{2}}$.
Lemma 8.
Consider the known-probability case. For the AIR-KP policy with the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(\frac{1}{2},1)$, there exists a constant $c_{10}$ independent of $T$ such that when $t\leq T-c_{10}$, under the good event, we have
  1. 1.
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=1$ and $Z_{j}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=0$.
  2. 2.
    $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{t,j^{t}}^{\mathcal{A}})=0$.
Lastly, we can prove the constant regret bound.
$\text{Reg}^{\mathcal{A}}(T)$ $\leq\sum_{t=1}^{T-c_{10}}r_{\phi}\mathbb{P}\left(\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{t,j^{t}}^{\mathcal{A}})>0\right)+c_{10}r_{\phi}$
$\leq\sum_{t=1}^{T-c_{10}}r_{\phi}\frac{c_{8}}{(T-t+1)^{2}}+c_{10}r_{\phi}\leq\left(\frac{\pi^{2}}{6}c_{8}+c_{10}\right)r_{\phi}.$
B.9 Proof of Lemma 6
In this case, by Hoeffding’s inequality, it holds that
$\mathbb{P}\left(|d_{j}^{t}-Z_{j}^{t}|\geq\sqrt{(T-T_{k-1}+1)\log(T-t+1)}\right)$
$=$ $\mathbb{P}\left(\left|p_{j}(T-T_{k-1}+1)-\sum_{\ell=T_{k-1}}^{T}Y_{j}^{\ell}\right|\geq\sqrt{(T-T_{k-1}+1)\log(T-t+1)}\right)$
$\leq$ $\frac{2}{(T-t+1)^{2}}.$
B.10 Proof of Lemma 7
First, by Hoeffding’s inequality, we have
$\mathbb{P}\left(d_{j}^{t}\leq(T-t+1)p_{j}-\sqrt{(t-T_{k-1})\log(T-t+1)}\right)$
$=$ $\mathbb{P}\left((t-T_{k-1})p_{j}-\sum_{\ell=T_{k-1}}^{t-1}Y^{\ell}_{j}\leq-\sqrt{(t-T_{k-1})\log(T-t+1)}\right)$
$\leq$ $\frac{1}{(T-t+1)^{2}}.$
Therefore, with probability greater than $1-\frac{1}{(T-t+1)^{2}}$, we have $d_{j}^{t}\geq(T-t+1)p_{j}-\sqrt{(t-T_{k-1})\log(T-t+1)}\geq(T-t+1)p_{j}-\sqrt{(T-t+1)^{1/\beta}\log(T-t+1)}$. Since $\frac{1}{2\beta}<1$, there exists a constant $c_{7}$ such that when $t\leq T-c_{7}$, $\boldsymbol{d}^{t}\geq\frac{T-t+1}{2}\boldsymbol{p}\geq\boldsymbol{2}$. Similar to Proposition 3, we can prove that $\boldsymbol{u}^{t}$ is an optimal solution to $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$ when $\boldsymbol{d}^{t}\geq\boldsymbol{2}$. ∎
B.11 Proof of Lemma 8
In the following, we always consider the good event. Consider the case when a type-$j$ customer arrives at time $t$ and the AIR-KP policy accepts his request. Similar to Proposition 4, given $t\in[T_{k-1},T_{k})$ and $t\leq T-c_{7}$, we have
$\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$ $\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\|\boldsymbol{d}^{t}-\boldsymbol{Z}^{t}\|_{\infty}\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
$\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\sqrt{2(T-t+1)^{1/\beta}\log(T-t+1)}.$
Since $\frac{1}{2\beta}<1$, there exists a constant $c_{9}$ such that when $t\leq T-c_{9}$, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=1$ and $Z_{j^{t}}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ otherwise. Then, we set $c_{10}=\max\{c_{7},c_{9},3\}$, and finish the proof. ∎
B.12 Proof of Theorem 4
We discuss different time intervals.
  1. 1.
    For the time interval $\left[1,\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-1}}\right\rceil\right)$, the proofs of Lemmas 6, 7 and 8 still hold, and hence the regret at time $t$ can be bounded by $\frac{c_{8}}{(T-t+1)^{2}}r_{\phi}$.
  2. 2.
    For the time interval $\left[\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-1}}\right\rceil+1,T\right]$, the regret is no greater than $T^{(1/2+\epsilon)\cdot\beta^{M-1}}r_{\phi}$.
To summarize, given the resolving time set $\mathcal{T}^{\mathcal{K}}$, the regret can be bounded as
$\text{Reg}^{\mathcal{A}}(T)$ $\leq\sum_{t=1}^{\left\lceil T-T^{(1/2+\epsilon)\cdot\beta^{M-1}}\right\rceil}\frac{c_{8}}{(T-t+1)^{2}}r_{\phi}+T^{(1/2+\epsilon)\cdot\beta^{M-1}}\cdot r_{\phi}$
$\leq\frac{c_{8}\pi^{2}}{6}r_{\phi}+T^{(1/2+\epsilon)\cdot\beta^{M-1}}r_{\phi}=\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-1}}).$
B.13 Proof of Lemma 3
For the known-probability problem, we need to prove similar results in Lemmas 7 and 8. Similar to Lemma 7, we have
$d_{j}^{t}\geq(T-t+1)p_{j}-\sqrt{(t-T_{k-1})\log(T-t+1)}\geq(T-t+1)p_{j}-\sqrt{\omega\log(T-t+1)},$
with probability $1-\frac{1}{(T-t+1)^{2}}$. There exists a constant $\tilde{\zeta}^{P}$ such that when $t\leq T-\tilde{\zeta}^{P}$, $\boldsymbol{d}^{t}\geq\frac{(T-t+1)}{2}\boldsymbol{p}\geq\boldsymbol{2}$. Moreover, $\tilde{\zeta}^{P}$ is the minimal positive integer $x$ satisfying $\frac{(x+1)\min_{j}p_{j}}{2}-\sqrt{\omega\log(x+1)}\geq 0$. Therefore, we can deduce that $\tilde{\zeta}^{P}=\mathcal{O}(\sqrt{\omega}\log\omega)=\tilde{\mathcal{O}}(\sqrt{\omega})$.
Similar to Lemma 8, under the good event, we have
$\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$ $\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
$\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\sqrt{(T-t+1+\omega)\log(T-t+1)}.$
Then, there exists a constant $\tilde{\theta}^{P}$ such that when $t\leq T-\tilde{\theta}^{P}$, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=1$ and $Z_{j}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=0$. Moreover, $\tilde{\theta}^{P}$ is the minimal integer $x$ satisfying $\frac{(x+1)\min_{j}p_{j}}{4}-\sqrt{(x+1+\omega)\log(x+1)}\geq 1$, implying that $\tilde{\theta}^{P}=\mathcal{O}(\sqrt{\omega}\log\omega)=\tilde{\mathcal{O}}(\sqrt{\omega})$. Then, the regret bound can be bounded as $\text{Reg}^{\mathcal{A}}(T)\leq\left(\frac{\pi^{2}}{6}c_{8}+\tilde{\theta}^{P}\right)r_{\phi}=\tilde{\mathcal{O}}(\sqrt{\omega})$. ∎
B.14 Proof of Proposition 5
For the unknown-probability case, we only need to prove similar results in Propositions 3 and 4.
Lemma 9.
For the resolving schedule $\mathcal{T}^{\mathcal{K},P}(\omega)$, there exist two constants $\eta^{P}=\mathcal{O}(\omega)$ and $\zeta^{P}=\mathcal{P}(\omega\log\omega)$ independent of $T$. When $t\in[\eta^{P},T-\zeta^{P}]$, with probability $1-\frac{c_{10}}{(T-t+1)^{2}}-\frac{c_{11}}{(\min\{T-t+1,t-1\})^{2}}$, we have
  1. 1.
    $d_{j}^{t}\geq\frac{p_{j}(T-t+1)}{2}\geq 2$ for any $j$.
  2. 2.
    $\boldsymbol{u}^{t}$ is an optimal solution to the surrogate LP $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$.
Lemma 10.
For the resolving schedule $\mathcal{T}^{\mathcal{K},P}(\omega)$, there exist two constants $\iota^{P}=\mathcal{O}(\omega)$ and $\theta^{P}=\mathcal{O}(\omega\log\omega)$ independent of $T$ such that when $t\in[\iota^{P},T-\theta^{P}]$, under the good event, we have
  1. 1.
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=1$ and $Z_{j}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$ if $x_{t,j^{t}}^{\mathcal{A}}=0$.
  2. 2.
    $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{t,j^{t}}^{\mathcal{A}})=0$.
Then, the regret bound can be bounded as $\text{Reg}^{\mathcal{A}}(T)\leq\left(\frac{\pi^{2}}{6}(c_{10}+c_{11})+\iota^{P}+\theta^{P}\right)r_{\phi}=\mathcal{O}(\omega)$.
Proof of Lemma 9. To derive a similar result in Proposition 3, we only need to show that $\boldsymbol{d}^{t}\geq\frac{T-t+1}{2}\boldsymbol{p}\geq\boldsymbol{2}$ with high probability. Consider the time period $t\in[T_{k-1},T_{k})$.
  1. 1.
    When $T_{k-1}<\lceil\frac{T}{2}\rceil$, we have
    $d_{j}^{t}$ $\geq(T-t+1)p_{j}-(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\sqrt{(t-T_{k-1})\log(T-t+1)}$
    $\geq(T-t+1)p_{j}-(T-t+\omega+1)\sqrt{\frac{\log(t-1)}{t-\omega-1}}-\sqrt{\omega\log(T-t+1)},$
    with probability $1-\frac{1}{(t-1)^{2}}-\frac{1}{(T-t+1)^{2}}$. There exist constants $\eta^{P}_{j}$ and $\zeta^{P}_{j}$ such that when $t\in[\eta^{P}_{j},\min\{T-\zeta^{P}_{j},\lceil\frac{T}{2}\rceil+\omega\}]$, we have $\sqrt{\frac{\log(t-1)}{t-\omega-1}}\leq\frac{p_{j}}{4}$, $\frac{p_{j}}{4}(T-t+1)-\frac{p_{j}\omega}{4}-\sqrt{\omega\log(T-t+1)}\geq 0$, and hence $d_{j}^{t}\geq\frac{p_{j}(T-t+1)}{2}\geq 2$. Moreover, we have $\eta^{P}_{j}=\mathcal{O}(\omega)$ and $\zeta_{j}^{P}=\mathcal{O}(\omega\log\omega)$.
  2. 2.
    When $T_{k-1}\geq\lceil\frac{T}{2}\rceil$, we have
    $d_{j}^{t}$ $\geq(T-t+1)p_{j}-\sqrt{2(T-T_{k-1}+1)\log(T-t+1)}-\sqrt{(t-T_{k-1})\log(T-t+1)}$
    $\geq(T-t+1)p_{j}-\sqrt{2(T-t+\omega+1)\log(T-t+1)}-\sqrt{\omega\log(T-t+1)}$
    $\geq(T-t+1)p_{j}-\sqrt{2(T-t+1)\log(T-t+1)}-\sqrt{2\omega\log(T-t+1)}-\sqrt{\omega\log(T-t+1)},$
    with probability $1-\frac{2}{(T-t+1)^{2}}$. Then, there exists a constant $\bar{\zeta}^{P}_{j}$ such that when $t\leq T-\bar{\zeta}^{P}_{j}$, we have $d_{j}^{t}\geq\frac{p_{j}(T-t+1)}{2}\geq 2$. Moreover, we have $\bar{\zeta}^{P}_{j}=\mathcal{O}(\omega\log\omega)$.
To summarize, by defining $\eta^{P}=\max_{j}\eta^{P}_{j}=\mathcal{O}(\omega)$ and $\zeta^{P}=\max\{\max_{j}\zeta^{P}_{j},\max_{j}\bar{\zeta}^{P}_{j}\}=\mathcal{O}(\omega\log\omega)$, we finish the proof.
Proof of Lemma 10. Consider the case when a type-$j$ customer arrives at time $t\in[T_{k-1},T_{k})$ and the AIR policy accepts his request. Similar to Proposition 4, we discuss two cases.
  1. 1.
    When $T_{k-1}<\lceil\frac{T}{2}\rceil$, we have
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\kappa_{4}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-t+1+\omega)\sqrt{\frac{\log(t-1)}{t-\omega-1}}-\kappa_{4}\sqrt{(T-t+\omega+1)\log(T-t+1)}.$
    There exists a constant $\iota^{P}_{j}$ such that when $t\geq\iota^{P}_{j}$, we have $2\kappa_{4}\sqrt{\frac{\log(t-1)}{t-\omega-1}}\leq\frac{p_{j}}{8}$. Moreover, there exists a constant $\theta^{P}_{j}$ such that when $t\leq T-\theta^{P}_{j}$, we have $\frac{(T-t+1)p_{j}}{8}-\frac{\kappa_{4}\omega p_{j}}{8}-\kappa_{4}\sqrt{(T-t+\omega+1)\log(T-t+1)}.$ Therefore, when $t\in[\iota^{P}_{j},\min\{\lceil\frac{T}{2}\rceil+\omega,T-\theta^{P}_{j}\}]$, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$. Moreover, we have $\iota_{j}^{P}=\mathcal{O}(\omega)$ and $\theta^{P}_{j}=\mathcal{O}(\omega\log\omega)$.
  2. 2.
    When $T_{k-1}\geq\lceil\frac{T}{2}\rceil$, we have
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(\sqrt{2}+1)\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $=$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(\sqrt{2}+1)\sqrt{(T-t+k+1)\log(T-t+1)}.$
    There exists a constant $\bar{\theta}^{P}_{j}$ such that when $\lceil\frac{T}{2}\rceil\leq t\leq T-\bar{\theta}^{P}_{j}$, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)\geq 1$. Moreover, we have $\bar{\theta}^{P}_{j}=\mathcal{O}(\omega\log\omega)$. Then, similar to Proposition 4, we can finish the proof.
B.15 Proof of Lemma 4
Similar to Lemma 3, we only need the following inequalities:
$d_{j}^{t}\geq(T-t+1)p_{j}-\sqrt{(t-T_{k-1})\log(T-t+1)}\geq(T-t+1)p_{j}-\sqrt{2(T-t+1)\log(T-t+1)}$
$\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$ $\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
$\geq\frac{p_{j}(T-t+1)}{4}-\kappa_{7}\sqrt{\left(2(T-t+1)+1\right)\log(T-t+1)}.$
Then, following a similar proof, we can derive the constant bound $\mathcal{O}(1)$. ∎
B.16 Proof of Proposition 6
Similar to Proposition 5, we only need the following inequalities:
  1. 1.
    When $T_{k-1}<\lceil\frac{T}{2}\rceil$, we have
    $d_{j}^{t}$ $\geq(T-t+1)p_{j}-(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\sqrt{(t-T_{k-1})\log(T-t+1)}$
    $\geq(T-t+1)p_{j}-\left(2(T-t+1)+1\right)\sqrt{\frac{2\log(t-1)}{t+1}}-\sqrt{\left(2(T-t+1)+1\right)\log(T-t+1)},$
    and
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\kappa_{4}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}\left(2(T-t+1)+1\right)\sqrt{\frac{2\log(t-1)}{t+1}}-\kappa_{4}\sqrt{\left(2(T-t+1)+1\right)\log(T-t+1)}.$
  2. 2.
    When $T_{k-1}\geq\lceil\frac{T}{2}\rceil$, we have
    $d_{j}^{t}$ $\geq(T-t+1)p_{j}-(T-T_{k-1}+1)\sqrt{\frac{\log(T-t+1)}{T_{k-1}-1}}-\sqrt{(t-T_{k-1})\log(T-t+1)}$
    $\geq(T-t+1)p_{j}-\left(2(T-t+1)+1\right)\sqrt{\frac{\log(T-t+1)}{\left(2(T-t+1)+1\right)}}-\sqrt{\left(2(T-t+1)+1\right)\log(T-t+1)},$
    and
    $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j)$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}(T-T_{k-1}+1)\sqrt{\frac{\log(t-1)}{T_{k-1}-1}}-\kappa_{4}\sqrt{(T-T_{k-1}+1)\log(T-t+1)}$
    $\geq$ $\frac{p_{j}(T-t+1)}{4}-\kappa_{4}\left(2(T-t+1)+1\right)\sqrt{\frac{\log(t-1)}{\left(2(T-t+1)+1\right)}}-\kappa_{4}\sqrt{\left(2(T-t+1)+1\right)\log(T-t+1)}.$
Then, following a similar proof, we can derive the constant bound $\mathcal{O}(1)$. ∎
Appendix C Numerical Details
In this section, we provide the omitted details in Section 4.
C.1 OLP Algorithms
Algorithm 3 Argmax with Frequent Resolving (AFR) Policy
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$ and $\boldsymbol{N}^{1}\leftarrow\boldsymbol{0}$.
for $t=1,2,3,\dots,T$ do
     Compute the empirical estimations $\hat{p}_{j}^{t}\leftarrow N_{j}^{t}/(t-1)$ for each $j$.
     Solve the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ and obtain its optimal solution $\boldsymbol{y}^{*}$.
     Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $y^{*}_{j}\geq(T-t+1)\hat{p}_{j}^{t}-y^{*}_{j}$ then $\triangleright$ Argmax between $y^{*}_{j}$ and $(T-t+1)\hat{p}_{j}^{t}-y^{*}_{j}$
         Accept the request.
         Set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$.
     else
         Reject the request.
     end if
end for
Algorithm 4 Adaptive Allocation (ADA) Policy
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$ and $\boldsymbol{N}^{1}\leftarrow\boldsymbol{0}$.
for $t=1,2,3,\dots,T$ do
     Compute the empirical estimations $\hat{p}_{j}^{t}\leftarrow N_{j}^{t}/(t-1)$ for each $j$.
     Solve the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ and obtain its optimal solution $\boldsymbol{y}^{*}$.
     Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ then
         Accept the request with probability $y_{j}^{*}/\left((T-t+1)\hat{p}_{j}^{t}\right)$. $\triangleright$ Probabilistic Allocation.
         If accepted, set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$.
     else
         Reject the request.
     end if
end for
Algorithm 5 Simple and Fast (SFA) Policy
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$ and $\boldsymbol{q}^{1}=\boldsymbol{0}$.
for $t=1,2,3,\dots,T$ do
     Observe arrival type $j$.
     Set $\tilde{x}^{t}\leftarrow 1$ if $r_{j}>\boldsymbol{A}_{j}\cdot\boldsymbol{q}^{t}$ and $\tilde{x}^{t}\leftarrow 0$ otherwise.
     Compute $\boldsymbol{q}^{t+1}\leftarrow\boldsymbol{q}^{t}+\frac{1}{\sqrt{t}}(\boldsymbol{A}_{j}\tilde{x}^{t}-\boldsymbol{\rho})$.
     Compute $\boldsymbol{q}^{t+1}\leftarrow\max\{\boldsymbol{q}^{t+1},\boldsymbol{0}\}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ then
         Accept the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$.
     end if
end for
Algorithm 6 Decoupling Learning and Decision (DLD) Policy
Input: $T_{e}=\lfloor T^{2/3}\rfloor$, $\alpha_{e}=T^{-1/3}$ and $\alpha_{p}=T^{-2/3}$.
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$, $\boldsymbol{q}^{1}_{D}\leftarrow\boldsymbol{0}$ and $\boldsymbol{q}^{1}_{L}\leftarrow\boldsymbol{0}$.
for $t=1,2,3,\dots,T_{e}$ do
     Observe arrival type $j$.
     Set $\tilde{x}^{t}_{D}\leftarrow 1$ if $r_{j}>\boldsymbol{A}_{j}\cdot\boldsymbol{q}^{t}_{D}$ and $\tilde{x}^{t}_{D}\leftarrow 0$ otherwise.
     Compute $\boldsymbol{q}^{t+1}_{D}\leftarrow\boldsymbol{q}^{t}_{D}+\alpha_{e}(\boldsymbol{A}_{j}\tilde{x}^{t}_{D}-\boldsymbol{\rho})$.
     Compute $\boldsymbol{q}^{t+1}_{D}\leftarrow\max\{\boldsymbol{q}^{t+1}_{D},\boldsymbol{0}\}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ then
         Accept the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$.
     end if
     Set $\tilde{x}^{t}_{L}\leftarrow 1$ if $r_{j}>\boldsymbol{A}_{j}\cdot\boldsymbol{q}^{t}_{L}$ and $\tilde{x}^{t}_{L}\leftarrow 0$ otherwise.
     Compute $\boldsymbol{q}^{t+1}_{L}\leftarrow\boldsymbol{q}^{t}_{L}+\frac{1}{t}(\boldsymbol{A}_{j}\tilde{x}^{t}_{L}-\boldsymbol{\rho})$.
     Compute $\boldsymbol{q}^{t+1}_{L}\leftarrow\max\{\boldsymbol{q}^{t+1}_{L},\boldsymbol{0}\}$.
end for
if $t=T_{e}+1$ then
     Set $\boldsymbol{q}^{t+1}_{D}\leftarrow\boldsymbol{q}^{t+1}_{L}$.
end if
for $t=T_{e}+1,T_{e}+2,\dots,T$ do
     Observe arrival type $j$.
     Set $\tilde{x}^{t}_{D}\leftarrow 1$ if $r_{j}>\boldsymbol{A}_{j}\cdot\boldsymbol{q}^{t}_{D}$ and $\tilde{x}^{t}_{D}\leftarrow 0$ otherwise.
     Compute $\boldsymbol{q}^{t+1}_{D}\leftarrow\boldsymbol{q}^{t}_{D}+\alpha_{p}(\boldsymbol{A}_{j}\tilde{x}^{t}_{D}-\boldsymbol{\rho})$.
     Compute $\boldsymbol{q}^{t+1}_{D}\leftarrow\max\{\boldsymbol{q}^{t+1}_{D},\boldsymbol{0}\}$.
     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ then
         Accept the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$.
     end if
end for
Algorithm 7 Budget-Updating Fast (BUF) Policy
Input: Time set $\mathcal{T}=\{T-\lceil\frac{T}{2^{k}}\rceil:k=1,2,\dots,\lceil\log_{2}T\rceil\}$.
Initialize $\boldsymbol{b}^{1}\leftarrow T\boldsymbol{\rho}$, $\boldsymbol{d}^{1}=\boldsymbol{\rho}$ and $\boldsymbol{q}^{1}=\boldsymbol{0}$.
for $t=1,2,3,\dots,T$ do
     Observe arrival type $j$.
     Set $\tilde{x}^{t}\leftarrow 1$ if $r_{j}>\boldsymbol{A}_{j}\cdot\boldsymbol{q}^{t}$ and $\tilde{x}^{t}\leftarrow 0$ otherwise.
     if $\boldsymbol{b}^{t}-\boldsymbol{A}_{j}\geq\boldsymbol{0}$ and $\tilde{x}^{t}=1$ then
         Accept the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}-\boldsymbol{A}_{j}$.
     else
         Reject the request and set $\boldsymbol{b}^{t+1}\leftarrow\boldsymbol{b}^{t}$.
     end if
     if $t+1\in\mathcal{T}$ then
         Set $l\leftarrow t+1$ and $\boldsymbol{d}^{t+1}=\frac{\boldsymbol{b}^{t+1}}{T-t}$.
     else
         Set $\boldsymbol{d}^{t+1}\leftarrow\boldsymbol{d}^{t}$
     end if
     Set $\boldsymbol{q}^{t+1}=\boldsymbol{q}^{t}+\frac{1}{t-l+2}(\boldsymbol{A}_{j}\tilde{x}^{t}-\boldsymbol{d}^{t+1})$
end for
C.2 Numerical Setup of the Multi-resource Case
In this section, we present the parameters randomly generated for the multi-resource case with $m=10$ and $n=2$.
$\boldsymbol{A}=\begin{bmatrix}0.226&0.146\\ 0.957&0.916\\ 0.005&0.876\\ 0.457&0.790\\ 0.285&0.960\\ 0.572&0.736\\ 0.701&0.206\\ 0.093&0.642\\ 0.903&0.923\\ 0.743&0.789\\ \end{bmatrix}\quad\boldsymbol{\rho}=\begin{bmatrix}0.128\\ 0.805\\ 0.770\\ 0.695\\ 0.844\\ 0.647\\ 0.181\\ 0.564\\ 0.812\\ 0.694\\ \end{bmatrix}\quad\boldsymbol{p}=\begin{bmatrix}0.121\\ 0.879\\ \end{bmatrix}\quad\boldsymbol{r}=\begin{bmatrix}0.689\\ 0.710\\ \end{bmatrix}$

### v3

@@
 In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound.
-In this work, we bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$.
+In this work, we study the case where the inputs are drawn from an unknown finite-support distribution, and bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$.
 Moreover, when we are allowed to solve LPs only $M$ times, we propose an algorithm that can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret.
@@
 Consider a decision maker with $m$ types of resources facing sequentially arriving customers over $T$ discrete time periods.
-We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{++}^{m}$ is given and denotes the vector of average inventories per period.
+We assume the initial inventory is $T\boldsymbol{\rho}$, where $\boldsymbol{\rho}=(\rho_{1},\dots,\rho_{m})\in\mathbb{R}_{+}^{m}$ is given and denotes the vector of average inventories per period.
 Each customer arriving at period $t$ is characterized by a reward $r_{t}\in\mathbb{R}_{+}$ and a consumption vector $\boldsymbol{A}_{t}\in\mathbb{R}_{+}^{m}$.
@@
 We focus on the asymptotic order of the regret as $T$ increases.
-Bray (2024) shows that the regret is at least $\mathcal{O}(\log T)$ when the demand distribution has infinite supports.
+Bray (2024) shows that the regret is at least $\Omega(\log T)$ when the demand distribution has infinite supports.
 In this paper, we focus on the case with a finite-support demand distribution.
@@
 2.1 Argmax with Infrequent Resolving (AIR) Policy
-We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy. The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
+We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy.
+We use $\boldsymbol{e}_{j}$ to denote a vector of zeros except 1 at the $j$-th entry.
+The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
 Algorithm 1 Argmax with Infrequent Resolving (AIR) Policy
@@
      Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
-     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
+     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
          Accept the request.
@@
 We adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera and Banerjee (2021) to make accept/reject decisions.
-At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand.
-Intuitively, $d_{j}^{t}-u_{t}^{j}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected.
+At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand.
+Intuitively, $d_{j}^{t}-u_{j}^{t}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected.
 In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
@@
      Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
-     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u^{j}_{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
+     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
          Accept the request.
@@
 Time Horizon Regret Computation Time (s) # of Resolvings
-AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA SFA* DLD* BUF* AIR
-2500 2.5 1.5 7.7 45.6 62.3 48.3 0.062 13.3 13.3 0.002 0.003 0.002 13
-5000 2.2 1.2 10.5 57.6 82.6 59.0 0.064 26.4 26.6 0.004 0.005 0.004 13
-7500 2.2 1.6 12.0 66.6 96.4 65.7 0.065 39.4 39.3 0.005 0.007 0.005 13
-10,000 2.2 1.4 13.2 74.4 109.7 72.5 0.066 54.4 54.2 0.007 0.009 0.007 13
-12,500 2.1 1.2 14.3 80.9 118.8 76.0 0.077 66.6 66.5 0.009 0.012 0.009 15
-15,000 2.2 1.3 15.3 86.8 128.1 79.7 0.086 73.3 73.2 0.011 0.015 0.011 15
-17,500 2.2 1.1 16.6 92.1 136.0 82.9 0.083 84.4 84.3 0.012 0.016 0.012 15
-20,000 2.1 1.0 17.4 97.0 141.6 85.9 0.084 103.2 103.1 0.014 0.019 0.013 15
-100,000 2.2 - [^4] The case is not studied due to the extremely long computation time. - 192.0 260.1 126.6 0.141 - - 0.080 0.105 0.078 15
-200,000 2.1 - - 260.2 330.9 151.6 0.204 - - 0.169 0.219 0.163 15
-300,000 2.1 - - 313.8 379.2 166.1 0.231 - - 0.220 0.291 0.213 15
+AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA SFA* DLD* BUF* AIR AFR ADA
+2500 2.5 1.5 7.7 45.6 62.3 48.3 0.062 13.3 13.3 0.002 0.003 0.002 13 2500 2500
+5000 2.2 1.2 10.5 57.6 82.6 59.0 0.064 26.4 26.6 0.004 0.005 0.004 13 5000 5000
+7500 2.2 1.6 12.0 66.6 96.4 65.7 0.065 39.4 39.3 0.005 0.007 0.005 13 7500 7500
+10,000 2.2 1.4 13.2 74.4 109.7 72.5 0.066 54.4 54.2 0.007 0.009 0.007 13 10,000 10,000
+12,500 2.1 1.2 14.3 80.9 118.8 76.0 0.077 66.6 66.5 0.009 0.012 0.009 15 12,500 12,500
+15,000 2.2 1.3 15.3 86.8 128.1 79.7 0.086 73.3 73.2 0.011 0.015 0.011 15 15,000 15,000
+17,500 2.2 1.1 16.6 92.1 136.0 82.9 0.083 84.4 84.3 0.012 0.016 0.012 15 17,500 17,500
+20,000 2.1 1.0 17.4 97.0 141.6 85.9 0.084 103.2 103.1 0.014 0.019 0.013 15 20,000 20,000
+100,000 2.2 - [^4] The case is not studied due to the extremely long computation time. - 192.0 260.1 126.6 0.141 - - 0.080 0.105 0.078 15 - -
+200,000 2.1 - - 260.2 330.9 151.6 0.204 - - 0.169 0.219 0.163 15 - -
+300,000 2.1 - - 313.8 379.2 166.1 0.231 - - 0.220 0.291 0.213 15 - -
 Table 3: Regret and Computation Time of Different Policies when $m=10$, $n=2$ and $\alpha=\beta=0.7$. Superscript * Marks LP-free Policies.

### v4

@@
 Abstract
-Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management and advertising.
+Online linear programming (OLP) has gained significant attention from both researchers and practitioners due to its extensive applications, such as online auction, network revenue management, order fulfillment and advertising.
 Existing OLP algorithms fall into two categories: LP-based algorithms and LP-free algorithms.
@@
 In contrast, LP-free algorithm only requires first-order computations but induces a worse performance, lacking a constant regret bound.
-In this work, we study the case where the inputs are drawn from an unknown finite-support distribution, and bridge the gap between these two extremes by proposing an algorithm that achieves a constant regret while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$.
-Moreover, when we are allowed to solve LPs only $M$ times, we propose an algorithm that can guarantee an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret.
-Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times.
+In this work, we bridge the gap between these two extremes by proposing a well-performing algorithm, that solves LPs at a few selected time points and conducts first-order computations at other time points.
+Specifically, for the case where the inputs are drawn from an unknown finite-support distribution, the proposed algorithm achieves a constant regret (even for the hard “degenerate” case) while solving LPs only $\mathcal{O}(\log\log T)$ times over the time horizon $T$.
+Moreover, when we are allowed to solve LPs only $M$ times, we design the corresponding schedule such that the proposed algorithm can guarantee a nearly $\mathcal{O}\left(T^{(1/2)^{M-1}}\right)$ regret.
+Our work highlights the value of resolving both at the beginning and the end of the selling horizon, and provides a novel framework to prove the performance guarantee of the proposed policy under different infrequent resolving schedules.
+Furthermore, when the arrival probabilities are known at the beginning, our algorithm can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times, and a nearly $\mathcal{O}\left(T^{(1/2)^{M}}\right)$ regret by solving LPs only $M$ times.
 Numerical experiments are conducted to demonstrate the efficiency of the proposed algorithms.
@@
 Online linear programming (OLP) is a classical problem in online decision making.
-In this problem, a decision maker manages multiple types of resources, such as airplane seats or hotel rooms, with limited inventory.
-Customers arrive sequentially, each requesting a combination of resources (e.g., multi-leg flights or multi-day room reservations) and offering a bid price.
+In this problem, a decision maker manages multiple types of resources, such as airplane seats or products, with limited inventory.
+Customers arrive sequentially, each requesting a combination of resources (e.g., multi-leg flights or bundled products) and offering a bid price.
 Upon observing an arriving customer’s request and bid price, the decision maker needs to make an irrevocable decision to accept or reject the request in order to maximize the total expected revenue without violating resource constraints.
-OLP has applications in various industries, including online auction (Buchbinder et al.
+OLP has applications in various areas, including online auction (Buchbinder et al.
 2007), advertisement (Mehta et al.
@@
 The time periods are indexed forward, starting from period $t=1$ and ending at period $t=T$.
-In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) [^1] Our results can be easily extended to the case with multiple actions.
-without knowing future information.
-Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) linear program:
+In each time period $t$, after observing the arriving customer characterized by $(r_{t},\boldsymbol{A}_{t})$, the decision maker must irrevocably decide whether to accept or reject the customer ($x_{t}\in\{0,1\}$) without knowing future information.
+Given full information $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$, we can solve the following (offline) integer linear program:
 $\max_{\boldsymbol{x}}\quad$ $\sum_{t=1}^{T}r_{t}x_{t}$
@@
 Such a problem is commonly referred to as online linear programming (OLP).
-In this paper, we assume the customers’ features $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d.
+In this paper, we assume that the customers’ features $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d.
 from an unknown distribution $\mathcal{P}$ with $n\in\mathbb{Z}_{+}$ supports.
-Such an assumption is widely adopted in the literature (e.g., Jasin 2015, Xie et al.
-2023 and Chen et al.
+Such an assumption is widely adopted in the online resource allocation literature (e.g., Jasin 2015, Xie et al.
+2023 and Gupta 2024).
+It is also motivated by practical considerations.
+In many industries including airline and e-commerce, the resource combinations and the prices are designed by the seller, and only customers interested in one of these options will be considered.
+Thus, the number of customer types is the same as the number of provided options, which is typically finite.
+Moreover, we assume that the time horizon $T$ is known in advance.
+For example, in the airline industry, the time horizon is determined by the known departure dates of the flights (see, e.g., Talluri and Van Ryzin 1998, Jasin 2015, Chen et al.
 2024).
-With this assumption, the OLP problem can be simplified as follows:
-There are $n$ types of arrivals with corresponding coefficient groups $\{(r_{j},\boldsymbol{A}_{j})\}_{j=1}^{n}$, and each arrival’s type is drawn according to the distribution $\{p_{j}\}_{j=1}^{n}$ (with $\sum_{j=1}^{n}p_{j}=1$).
+Without the knowledge of the time horizon, it will be hard to derive a policy with a good performance guarantee because the remaining time plays a crucial role in the decision process (see, e.g., Agrawal et al.
+2014, Jasin and Kumar 2012 and Bumpensanti and Wang 2020).
+With the above assumptions, the OLP problem can be simplified as follows:
+There are $n$ types of customers with type-$j$ customers characterized by the coefficients $(r_{j},\boldsymbol{A}_{j})$, and each arriving customer’s type is drawn according to a distribution $\{p_{j}\}_{j=1}^{n}$ (with $\sum_{j=1}^{n}p_{j}=1$).
 Slightly abusing the notations, we define $\boldsymbol{r}\in\mathbb{R}_{+}^{n}$ and $\boldsymbol{A}\in\mathbb{R}_{+}^{m\times n}$ as the reward vector and the consumption matrix for all types of customers, respectively.
@@
 where $\Pi$ is the set of all non-anticipative policies, $x^{t}_{\mu,j}$ denotes whether the policy $\mu$ accepts a type-$j$ customer at time $t$, and $Y_{j}^{t}\in\{0,1\}$ is the random indicator whether a type-$j$ customer arrives at time $t$.
-If we relax the integer constraints in (2) and replace all random variables with their expectations, then the model is referred to as the “fluid model”.
 Solving the optimal policy $\mu^{*}$ for (2) is computationally intractable even if the underlying distribution $\mathcal{P}$ is known.
@@
 Furthermore, a lower-order regret typically implies a better performance, and designing a policy with constant regret, i.e., $\text{Reg}^{\mu}(T)=\mathcal{O}(1)$, is highly desirable.
-Broadly speaking, OLP algorithms in prior studies can be categorized into two types: LP-based algorithms and LP-free algorithms.
-LP-based algorithms typically make decisions based on the optimal primal/dual solution to the fluid LP in each period, offering strong performance guarantees but requiring solving a large number of LPs (e.g., Li and Ye 2022, Xie et al.
+Since (2) is computationally intractable, researchers typically design heuristic policies based on a relaxed problem.
+Specifically, if we relax the integer constraints in (2) and replace all random variables with their (estimated) expectations, then we can derive the following linear program (let $y_{j}$ denote the number of accepted type-$j$ customers, and $\hat{\boldsymbol{p}}^{t}$ denote the estimated probabilities at period $t$)
+$\begin{split}\bar{V}(T):=\max_{\boldsymbol{y}\geq\boldsymbol{0}}\quad&\boldsymbol{r}^{\mathsf{T}}\boldsymbol{y}\\ \text{s.t.}\quad&\boldsymbol{A}\boldsymbol{y}\leq T\boldsymbol{\rho},\\ &\boldsymbol{y}\leq\hat{\boldsymbol{p}}^{1}\cdot T,\ \forall j,\end{split}$ (4)
+which is referred to as the “fluid model” (at period 1).
+Given a period $t$, some algorithms may update the remaining inventory and the estimated future arrivals in (4), and resolve the (updated) fluid linear program.
+Broadly speaking, based on the number of LP resolvings, OLP algorithms in prior studies can be categorized into two types: LP-based algorithms and LP-free algorithms.
+LP-based algorithms typically make decisions based on the optimal primal/dual solution to the updated fluid LP (4) in each period, offering strong performance guarantees but requiring solving a large number of LPs (e.g., Li and Ye 2022, Xie et al.
 2023, Chen et al.
@@
 Although commercial solvers can efficiently solve LPs, frequent LP resolving remains computationally expensive especially for large-scale or time-sensitive problems.
+For example, the leading hotel booking platform Booking.com received about 560 million visits per month during 2022-2024, implying about 5 milliseconds average inter-arrival time (Statista 2024).
+In contrast, it typically consumes from seconds to minutes to solve linear programs of practical problems (see, e.g., Mittelmann 2024).
+Moreover, the optimal basis may change during the time horizon, and hence for each period we cannot simply use the last-period optimal solution to significantly speed up LP solving.
 In response, researchers recently propose LP-free algorithms that typically use gradient descent methods to derive an approximate solution converging to the optimal dual solution to the fluid LP, and makes decisions based on the approximate solution.
@@
 Specifically, we propose an algorithm which achieves $\mathcal{O}(1)$ regret for the OLP problem while requiring solving LPs only $\mathcal{O}(\log\log T)$ times.
-In our algorithm, we concentrate the resolving (of an approximate LP) periods 1) on the beginning of the time horizon and 2) toward the end of the time horizon.
-The first set of resolvings aim to update the policy when the data is scarce so that correcting learning errors is important, while the latter set of resolvings aim to update the policy when the inventory is running out.
-Between two resolvings, we use the solution of an approximate LP to guide the allocation.
-Moreover, when we are allowed to solve LPs for up to $M$ times, we propose an algorithm that can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret.
+In our algorithm, we concentrate the resolving (of the fluid LP) periods 1) on the beginning of the time horizon and 2) toward the end of the time horizon.
+The first set of resolvings aim to update the policy when the data is scarce so that correcting learning errors is important, while the latter set of resolvings aim to update the policy when the inventory is running out so that the optimal policy varies dramastically.
+Between two resolvings, we use the optimal fluid solution of the latest resolving plus some first-order computations to guide the allocation.
+Moreover, when we are allowed to solve LPs for up to $M$ times, we propose an algorithm that can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M-1}}\right)$ regret, where $\epsilon$ can be any positive constant.
+If we take $\epsilon\to 0$, the regret bound is about $\mathcal{O}\left(T^{(1/2)^{M-1}}\right)$.
 In addition, we show that the proposed algorithms can be easily adapted to the settings where the arrival probabilities of each type of request are known, which also lead to near-optimal regrets in those settings.
 In this case, a modified algorithm can achieve an $\mathcal{O}\left(T^{(1/2+\epsilon)^{M}}\right)$ regret by solving LPs only $M$ times.
-Therefore, our results demonstrate that one can achieve near optimal regrets only with a few resolvings, and depict the precise tradeoff between the frequency of resolving and the performance of the algorithm.
-We would like to highlight the technical contribution of this work.
-In the literature of LP-based algorithms for solving OLP, the algorithm usually solves an updated fluid model in each time period (e.g., Vera and Banerjee 2021, Xie et al.
+Therefore, our results demonstrate that one can achieve near optimal regrets only with a few resolvings, and depict the precise tradeoff between the frequency of resolving and the performance of the proposed algorithm.
+In addition to the strong performance of our proposed algorithm, we would like to highlight one especially significant feature of our algorithm.
+In many previous studies, an important “non-degeneracy” assumption is imposed on the underlying input data and such an assumption would greatly affect the performance of proposed algorithms.
+Particularly, the “non-degeneracy” assumption refers to the assumption that the fluid model (4) with the estimation $\hat{\boldsymbol{p}}^{1}$ replaced by the true value $\boldsymbol{p}$ (called the “no-learning fluid model”) is non-degenerate.
+In some works (e.g., Gupta 2024, Wei et al.
+2023), a $\delta$ is defined to be a measure of the distance between the current inventory configuration $T\boldsymbol{\rho}$ and the nearest inventory configuration under which the no-learning fluid model is degenerate, and the regret result is dependent on $1/\delta$ (thus their regret bounds dramastically increase and tend to infinity as $\delta$ tends to zero).
+Importantly, our results do not rely on the non-degeneracy assumption.
+In fact, to the best of our knowledge, our algorithm is the first that achieves a constant regret with such few resolvings for the case without the non-degeneracy assumption and distribution knowledge.
+To distinguish different results, we use $\mathcal{O}_{\delta}(\cdot)$ ($\mathcal{O}(\cdot)$, resp.) to denote regret bounds containing (without, resp.) $1/\delta$.
+Finally, we would like to highlight the technical contribution of this work.
+In the literature of LP-based algorithms for solving OLP, the algorithm usually solves an updated fluid model in each time period (e.g., Vera and Banerjee 2020, Xie et al.
 2023, Chen et al.
 2024).
-This approach allows bridging their policies with the optimal hindsight policy through two LPs: the fluid model and the hindsight benchmark.
-In contrast, our approach solves the updated fluid model only at a few selected periods ($\mathcal{O}(\log\log T)$ periods), and hence cannot access the optimal solution of the updated fluid model in most periods, posing a challenge for the analysis.
-In order to overcome this technical challenge, we approximate the optimal solution of the updated fluid model based on the latest obtained fluid solution and some first-order computations.
-While this solution is sub-optimal for the updated fluid model, we prove that it is optimal to a surrogate LP with high probability.
+This approach allows updating the probability estimation frequently and bridging their policies with the optimal hindsight policy through two LPs: the fluid model and the hindsight benchmark.
+In contrast, our approach solves the updated fluid model only at a few selected periods ($\mathcal{O}(\log\log T)$ periods), and hence cannot update the probability estimation and access the optimal solution of the updated fluid model in most periods, posing challenges for the analysis.
+In order to overcome these technical challenges, we approximate the optimal solution of the updated fluid model based on the latest obtained fluid solution and some first-order computations.
+While this solution is sub-optimal for the updated fluid model for the corresponding period, we prove that it is optimal to a surrogate LP with high probability.
 Therefore, we can bridge our policy with the optimal hindsight policy through the surrogate LP and the hindsight LP, from which we can obtain the desired result.
@@
 Unknown Distribution.
-We start with the stream assuming unknown arrival probabilities.
+We start with the stream assuming unknown type distribution.
 Our work closely relates to studies under the random input assumption, where the coefficients $\left\{(r_{t},\boldsymbol{A}_{t})\right\}_{t=1}^{T}$ are generated i.i.d.
 from an unknown distribution $\mathcal{P}$.
-We focus on the asymptotic order of the regret as $T$ increases.
-Bray (2024) shows that the regret is at least $\Omega(\log T)$ when the demand distribution has infinite supports.
-In this paper, we focus on the case with a finite-support demand distribution.
+For the underlying distribution, there are two diverging assumptions: finite-support distribution and continuous-support distribution.
+We start with the finite-support distribution assumption, which is the same as our setting.
+Under such an assumption, the distribution $\mathcal{P}$ is supported by finite bounded points such that arrivals can be categorized into finite types.
 We first review the so called LP-based algorithms which require solving many LPs but guarantee good performance.
-For example, under certain non-degeneracy assumptions, Jasin (2015) proposes an $\mathcal{O}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times.
+For example, under the non-degeneracy assumption (i.e., $\delta>0$), Jasin (2015) proposes an $\mathcal{O}_{\delta}(\log^{2}T)$-regret algorithm that requires solving LPs $\mathcal{O}(\log T)$ times.
 Subsequently, Chen et al.
-(2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise.
-Under the non-degeneracy assumption and some distribution assumptions, Li and Ye (2022) provide an $\mathcal{O}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times.
-Then, under similar assumptions, Ma et al.
-(2024) provide an $\mathcal{O}(\log T)$-regret algorithm requiring solving LPs $\mathcal{O}(T)$ times.
-Recently, Xie et al.
+(2024) consider a similar algorithm that requires solving LPs $T$ times, achieving $\mathcal{O}_{\delta}(1)$ regret under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T}\log T)$ otherwise.
+Wei et al.
+(2023) propose a primal-dual algorithm that solves LPs $T$ times and the corresponding regret is $\mathcal{O}_{\delta}(1)$ under the non-degeneracy assumption and $\mathcal{O}(\sqrt{T})$ in general.
+Xie et al.
 (2023) remove the non-degeneracy assumption and provide an OLP algorithm with $\mathcal{O}(1)$ regret but still requiring solving LPs $T$ times.
-Moreover, there are studies on infrequent resolving algorithms under additional assumptions.
-Given a complete sample path (i.e., $T$ samples), Freund and Banerjee (2019) propose an algorithm with lazy resolving (i.e., $\mathcal{O}(\log\log T)$ expected LP solving times) and prove a constant regret.
-Given a starting point with an accuracy of $\mathcal{O}(1/T)$, Ma et al.
-(2024) propose an infrequent resolving algorithm that can achieve $\mathcal{O}(\log T)$ regret under the non-degeneracy assumption by solving LPs $O(\log T)$ times.
-Both assumptions are stringent when $T$ is large.
-In this work, without such assumptions, we propose an algorithm that can guarantee a constant regret by solving LPs $\mathcal{O}(\log\log T)$ times.
-Table 1: Comparison among OLP Algorithms (i.e., Unknown-Probability Case).
-Paper
-Regret
-(Non-degenerate Case)
-Regret
-(Degenerate Case)
-# of Resolvings
-Jasin (2015) $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log T)$
-Li and Ye (2022) $\mathcal{O}(\log T\log\log T)$ $\mathcal{O}(T)$
-Chen et al. (2024) $\mathcal{O}(1)$ $\mathcal{O}(\sqrt{T}\log T)$ $\mathcal{O}(T)$
-Xie et al. (2023) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
-Ma et al. (2024) $\mathcal{O}(\log T)$ $\mathcal{O}(T)$
-Freund and Banerjee (2019)
-(Requiring $T$ samples )
-$\mathcal{O}(1)$ $\mathcal{O}(1)$
-$\mathcal{O}(\log\log T)$
-(In expectation )
-Ma et al. (2024)
-(Requiring $\mathcal{O}(1/T)$-accuracy starting point )
-$\mathcal{O}(\log T)$ $\mathcal{O}(\log T)$
-Li et al. (2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00
-Gao et al. (2024) $\mathcal{O}(T^{1/3})$ 00
-Ma et al. (2024) $\mathcal{O}(\log^{2}T)$ 00
-This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
-This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $M$
-Although LP-based algorithms have good performance guarantees, frequent LP solving can be computational expensive for large-scale or time-sensitive problems.
-To address these computational challenges, recent studies have developed LP-free algorithms that never solve full LPs.
-Most of these studies analyze the regret in absence of budget constraints and derive bounds for both the regret and the constraint violation, i.e., the cumulative over-budget consumption.
-[^2] As Li et al.
-(2020) note, although an LP-free algorithm can be modified to a feasible policy satisfying the budget constraints, deriving a regret bound for such a modified policy remains an open question.
-The pioneering work Li et al.
-(2020) adopts a stochastic sub-gradient descent method to learn the optimal dual prices and provides an $\mathcal{O}(\sqrt{T})$-regret LP-free algorithm with $\mathcal{O}(\sqrt{T})$ constraint violation.
-This algorithm only requires first-order computations and never solves any full LP.
-Subsequently, researchers have studied variants of this algorithm and derived tighter bounds under the non-degeneracy assumption and the distribution assumptions in Li and Ye (2022).
+As mentioned, although LP-based algorithms have good performance guarantees, frequent LP solving can be computational expensive for large-scale or time-sensitive problems.
+To address these computational challenges, recent studies have developed LP-free algorithms without LP resolving.
+The pioneering works Balseiro et al.
+(2020) and Li et al.
+(2020) adopt stochastic gradient descent methods to learn the optimal dual prices and provide $\mathcal{O}(\sqrt{T})$-regret LP-free algorithms.
+[^1] The algorithm in Li et al.
+(2020) also allows consuming $\mathcal{O}(\sqrt{T})$ additional budget.
+As Li et al.
+(2020) note, although the proposed LP-free algorithm can be modified to a feasible policy satisfying the budget constraints, deriving a regret bound for such a modified policy remains an open question.
+These algorithms only require first-order computations and never solve any full LP.
+In this work, we achieve a balance between computational efficiency and algorithm performance by proposing an algorithm that achieves a constant regret bound by solving LPs $\mathcal{O}(\log\log T)$ times.
+Moreover, we also provide regret bounds when the number of resolvings is finite, which to the best of our knowledge has not been provided in previous literature.
+Table 1: Comparison among OLP Algorithms for Unknown-Probability Case.
+Paper Regret (Non-degenerate Case) Regret (Degenerate Case) # of Resolvings Distribution Assumption Jasin (2015) $\mathcal{O}_{\delta}(\log^{2}T)$ $\mathcal{O}(\log T)$ Finite Chen et al.
+(2024) $\mathcal{O}_{\delta}(1)$ $\mathcal{O}(\sqrt{T}\log T)$ $\mathcal{O}(T)$ Finite Wei et al.
+(2023) $\mathcal{O}_{\delta}(1)$ $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(T)$ Finite Xie et al.
+(2023) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$ Finite Li and Ye (2022) $\mathcal{O}_{\delta}(\log T\log\log T)$ $\mathcal{O}(T)$ Continuous Bray (2024) $\mathcal{O}_{\delta}(\log T)$ $\mathcal{O}(T)$ Continuous Ma et al.
+(2024) $\mathcal{O}_{\delta}(\log T)$ $\mathcal{O}(T)$ Continuous Balseiro et al.
+(2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00 General Li et al.
+(2020) $\mathcal{O}(\sqrt{T})$ $\mathcal{O}(\sqrt{T})$ 00 General Gao et al.
+(2024) $\mathcal{O}_{\delta}(T^{1/3})$ 00 Continuous Ma et al.
+(2024) $\mathcal{O}_{\delta}(\log^{2}T)$ 00 Continuous This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ Finite This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M-1}})$ $M$ Finite
+Then, we review papers with the continuous-support distribution assumption.
+In this case, the distribution $\mathcal{P}$ is assumed to be supported by a bounded and continuous set and the probability density function is both lower bounded (away from 0) and upper bounded.
+(Note that finite-support distributions do not satisfy the assumption due to the existence of mass points.
+) We still start with LP-based algorithms.
+For example, under the non-degeneracy assumption, Li and Ye (2022) provide an $\mathcal{O}_{\delta}(\log T\log\log T)$-regret algorithm from the dual perspective, and the algorithm requires solving LPs $T$ times.
+Then, Bray (2024) shows that the best possible regret bound for this problem is $\Omega(\log T)$.
+In addition, Bray (2024) and Ma et al.
+(2024) prove that the regret bound of Li and Ye (2022)’s algorithm is $\mathcal{O}_{\delta}(\log T)$ under the non-degeneracy assumption.
+Then, we introduce LP-free algorithms for this setting.
+The $\mathcal{O}(\sqrt{T})$ regret bounds established in Balseiro et al.
+(2020) and Li et al.
+(2020) also hold in this case.
+Subsequently, researchers have studied variants of this algorithm and tried to derive tighter bounds under the non-degeneracy assumption.
 For example, Gao et al.
-(2024) propose variants that improve the bounds for both regret and the constraint violation to $\mathcal{O}(T^{1/3})$.
-Similarly, Ma et al.
-(2024) propose an LP-free algorithm which guarantees an $\mathcal{O}(\log^{2}T)$ regret without constraint violation.
-In this work, we achieve a balance between computational efficiency and algorithm performance by proposing an algorithm that achieves a constant regret bound by solving LPs $\mathcal{O}(\log\log T)$ times.
-A summary of the existing results and our result is presented in Table 1.
-Another stream of research studies the OLP problem under the random permutation assumption.
+(2024) propose variants that improve the bounds to $\mathcal{O}_{\delta}(T^{1/3})$.
+[^2] The algorithm in Gao et al.
+(2024) also allows consuming $\mathcal{O}(T^{1/3})$ additional budget.
+Then, Ma et al.
+(2024) propose an LP-free algorithm which guarantees an $\mathcal{O}_{\delta}(\log^{2}T)$ regret.
+A summary of the existing results and our result is presented in Table 1 (where $\epsilon$ can be any positive constant).
+As mentioned, the regret bounds of some works are infinite under the degenerate case (i.e., $\delta=0$).
+We will use ‘-’ to denote such results.
+Moreover, there are some works extending the previous settings.
+For example, for any bounded distribution (subsuming previous two classes), Balseiro et al.
+(2020), Li et al.
+(2020), Gao et al.
+(2023) and Balseiro et al.
+(2022) provide LP-free algorithms with $\mathcal{O}(\sqrt{T})$ regrets.
+Jiang et al.
+(2020) and Balseiro et al.
+(2023) consider the case when the underlying distribution is non-stationary.
+Besbes and Zeevi (2012) and Ferreira et al.
+(2018) consider the pricing problem without the knowledge of the demand function.
+We also note another stream of research studies under the random permutation assumption.
 Here, the set $\{(r_{t},\boldsymbol{A}_{t})\}_{t=1}^{T}$ is adversarially chosen, but the arrival order is uniformly distributed over all the permutations.
@@
 We also highlight that Agrawal et al.
-(2014), Molinaro and Ravi (2014) and Gupta and Molinaro (2014) solve LPs $\mathcal{O}(\log T)$ times and their resolving periods concentrate at the beginning, sharing similarities with our resolving schedule.
+(2014), Molinaro and Ravi (2014) and Gupta and Molinaro (2014) solve LPs $\mathcal{O}(\log T)$ times and their resolving periods concentrate at the beginning, sharing similarities with the first half of our resolving schedule.
 Known distribution.
@@
 The former one studies the dynamic pricing problem under resource constraints (see Gallego and Van Ryzin 1994); the latter one studies the dynamic resource allocation problem (see Talluri and Van Ryzin 1998).
-Our work is closely related to the quantity-based NRM, in which most algorithms are LP-based.
+Our work is closely related to the quantity-based NRM, in which most works adopt the finite-support assumption (see Jiang et al.
+2022 and Besbes et al.
+2024 for some exceptions).
+Most early works only solve the fluid LP at the beginning and then design a policy based on the optimal solution.
 For example, Talluri and Van Ryzin (1998) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the bid-price control (BPC) policy, which uses the optimal dual variable to the fluid model as the values of resources and accepts a request if and only if the offered price is larger than the total value of resources.
@@
 Reiman and Wang (2008) prove an $\mathcal{O}(\sqrt{T})$ regret bound for the probabilistic allocation control (PAC) policy, which probabilistically accepts a request according to the ratio of the fluid model solution to the expected demand.
-Table 2: Comparison among Algorithms for Known-Probability Case.
-Paper
-Regret
-(Non-degenerate)
-Regret
-# of Resolvings
-Jasin and Kumar (2012) $\mathcal{O}(1)$ [^3] Bumpensanti and Wang (2020) prove that the regret of the algorithm in Jasin and Kumar (2012) is $\Theta(\sqrt{T})$. $\mathcal{O}(\log T)$
-Bumpensanti and Wang (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
-Arlotto and Gurvich (2019)
-(Single constraint )
-$\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
-Vera and Banerjee (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
-Vera et al. (2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$
-Sun et al. (2020) $\mathcal{O}(T^{3/8}(\log T)^{5/4})$ 0
-Reiman and Wang (2008) $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ 2
-Jasin and Kumar (2012) $\mathcal{O}(T^{\frac{1}{2^{M}}})$ $M$
-Bumpensanti and Wang (2020) $\mathcal{O}(T^{5/12})$ $\mathcal{O}(T^{5/12})$ 2
-Gupta (2024) $\mathcal{O}(1)$ 1
-This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$
-This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $M$
-Observing the potential to resolve the fluid model to reduce the regret, many works consider algorithms with resolving.
-For example, Cooper (2002) provides an example where frequent resolving may hurt the revenue.
-Then, considering the asymptotic regime, Reiman and Wang (2008) prove that a single resolving can reduce the regret of the PAC policy from $\Theta(\sqrt{T})$ to $o(\sqrt{T})$.
-Jasin and Kumar (2012) and Jasin and Kumar (2013) show that the PAC policy can significantly benefit from resolving, reducing its regret from $\Theta(\sqrt{T})$ to $\mathcal{O}(1)$, while neither BPC nor BLC can benefit.
-Specifically, under the non-degeneracy assumption, the PAC policy with periodic ($\mathcal{O}(T)$ times) or midpoint ($\mathcal{O}(\log T)$ times) resolving can achieve $\mathcal{O}(1)$ regret.
+Table 2: Comparison among Algorithms for Known-Probability Case (Finite-Support Distribution).
+Paper Regret (Non-degenerate) Regret (Degenerate) # of Resolvings Jasin and Kumar (2012) $\mathcal{O}_{\delta}(1)$ [^3] Bumpensanti and Wang (2020) prove that the regret of the algorithm in Jasin and Kumar (2012) is $\Theta(\sqrt{T})$.
+$\mathcal{O}(\log T)$ Bumpensanti and Wang (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ Vera and Banerjee (2020) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$ Vera et al.
+(2021) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(T)$ Banerjee and Freund (2024) $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ (In Expectation) Reiman and Wang (2008) $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ 2 Jasin and Kumar (2012) $\mathcal{O}_{\delta}(T^{\frac{1}{2^{M}}})$ $M$ Bumpensanti and Wang (2020) $\mathcal{O}(T^{5/12})$ $\mathcal{O}(T^{5/12})$ 2 Sun et al.
+(2020) $\mathcal{O}_{\delta}(T^{3/8}(\log T)^{5/4})$ 0 Gupta (2024) $\mathcal{O}_{\delta}(1)$ 1 This paper $\mathcal{O}(1)$ $\mathcal{O}(1)$ $\mathcal{O}(\log\log T)$ This paper $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ $M$
+Observing the potential to resolve the fluid model to reduce the regret, many works consider LP-based algorithms with more resolvings.
+For example, Reiman and Wang (2008) prove that a single resolving can reduce the asymptotic regret of the PAC policy from $\Theta(\sqrt{T})$ to $o(\sqrt{T})$.
+Jasin and Kumar (2012) and Jasin and Kumar (2013) show that the PAC policy can significantly benefit from resolving, while neither BPC nor BLC can benefit.
+Specifically, under the non-degeneracy assumption, the PAC policy with periodic ($\mathcal{O}(T)$ times) or midpoint ($\mathcal{O}(\log T)$ times) resolving can achieve $\mathcal{O}_{\delta}(1)$ regret.
 Without the non-degeneracy assumption, Bumpensanti and Wang (2020) show that the above regret in general is $\Omega(\sqrt{T})$, and provide a modified PAC policy with infrequent (i.e., $\mathcal{O}(\log\log T)$ times) resolving that guarantees $\mathcal{O}(1)$ regret.
+Note that the resolving schedule in Bumpensanti and Wang (2020) shares some similarities with the second half of our schedule, but their proof framework cannot deal with the case without distribution knowledge.
 Recently, Arlotto and Gurvich (2019) study an alternative interpretation of the fluid solution, which accepts the request if and only if the ratio of the primal solution to the expected demand is no less than $1/2$.
@@
 In our work, we refer to such policy as the “argmax policy” because it takes the action (accept/reject) with the larger value in the primal solution.
-Vera and Banerjee (2021) and Vera et al.
+Then, Vera and Banerjee (2020) and Vera et al.
 (2021) generalize this idea to the multi-constraint problem, but their policy requires solving LPs in every period.
-In our work, we also use the argmax policy as the base policy.
-We show that such a policy can guarantee $\mathcal{O}(1)$ regret by solving only $\mathcal{O}(\log\log T)$ LPs.
-In the LP-free regime, Sun et al.
-(2020) propose an LP-free algorithm with $\mathcal{O}(T^{3/8}(\log T)^{5/4})$ regret.
-Finally, we would like to mention some results on the case when the decision maker is only allowed to solve a finite number (i.e., $M$) of LPs under the NRM model.
+Subsequently, Banerjee and Freund (2024) propose a constant-regret algorithm whose expected number of resolvings is $\mathcal{O}(\log\log T)$.
+Similarly, the proof in Banerjee and Freund (2024) cannot handle the unknown-probability case under infrequent resolving.
+In our work, inspired by the argmax policy, we propose a constant-regret policy whose resolving schedule can be determined at the beginning and the number of resolvings is $\mathcal{O}(\log\log T)$.
+Then, we would like to mention some results on the case when the decision maker is only allowed to solve a finite number (i.e., $M$) of LPs under the NRM model.
 For this problem, Reiman and Wang (2008) show that the regret can be reduced to $\mathcal{O}(T^{\frac{1}{4}+\epsilon})$ if we can solve LPs twice.
 However, the proof cannot be directly extended to the multiple-resolving case.
-Under the non-degeneracy assumption, Jasin and Kumar (2012) prove that solving LP at $M$ periods can induce an $\mathcal{O}(T^{\frac{1}{2^{M}}})$ regret bound.
+Under the non-degeneracy assumption, Jasin and Kumar (2012) prove that solving LP at $M$ periods can induce an $\mathcal{O}_{\delta}(T^{\frac{1}{2^{M}}})$ regret bound.
 Due to the restricted choice of resolving times, Bumpensanti and Wang (2020) prove that the regret bound of their policy is $\mathcal{O}(T^{5/12})$ given $M=2$, which is worse than previous results.
-In addition, Gupta (2024) recently proposes an algorithm which only solves LP once at the beginning and then greedily makes decisions to minimize a sum-of-squares function in each period, whose regret is proved to be $\mathcal{O}(1)$ under the non-degeneracy assumption.
+Sun et al.
+(2020) propose an LP-free algorithm with $\mathcal{O}_{\delta}(T^{3/8}(\log T)^{5/4})$ regret under the non-degeneracy assumption.
+In addition, Gupta (2024) recently proposes an algorithm which only solves LP once at the beginning (to find the optimal basis) and then greedily makes decisions to minimize some function in each period, whose regret is proved to be $\mathcal{O}_{\delta}(1)$ under the non-degeneracy assumption.
 Compared to the above literature, our results can meet the single-resolving result of Reiman and Wang (2008) and extend it to the finite-resolving case.
@@
 We now introduce our policy in Algorithm 1, referred to as the Argmax with Infrequent Resolving (AIR) policy.
-We use $\boldsymbol{e}_{j}$ to denote a vector of zeros except 1 at the $j$-th entry.
+We use $\boldsymbol{e}_{j}$ to denote a vector of zeros except 1 at the $j$-th entry, and $[n]$ to denote the set $\{1,2,3,\dots,n\}$.
 The time set $\mathcal{T}$ in Algorithm 1 will be specified shortly.
@@
      Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
-     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
+     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$, $u_{j}^{t}>1$, and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
          Accept the request.
@@
 The algorithm begins with a predetermined time set $\mathcal{T}$ specifying the time points where resolving is needed.
-At each time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ as an approximation to the future demand and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$.
-The optimal solution $\boldsymbol{y}^{*}$ of the fluid LP provides an approximation to the optimal policy.
-At each time period $t\notin\mathcal{T}$, the decision maker updates an approximation $\boldsymbol{u}^{t}$ to the optimal solution in a “greedy” fashion as described in the algorithm.
-We adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera and Banerjee (2021) to make accept/reject decisions.
-At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and if $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$, where $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand.
-Intuitively, $d_{j}^{t}-u_{j}^{t}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if more should be accepted than rejected.
+At each resolving time point $t\in\mathcal{T}$, the decision maker updates the empirical estimations $\hat{\boldsymbol{p}}^{t}$ and solves the corresponding fluid LP based on the current inventory $\boldsymbol{b}^{t}$ and the estimation $\hat{\boldsymbol{p}}^{t}$.
+To implement our policy, we need to maintain two approximations in our algorithm: The vector $\boldsymbol{u}^{t}=\boldsymbol{y}^{*}$ approximates the numbers of accepted future customers (of different types) under the optimal policy, and the vector $\boldsymbol{d}^{t}=(T-t+1)\hat{\boldsymbol{p}}$ approximates the numbers of future customer arrivals (of different types).
+However, at each non-resolving time period $t\notin\mathcal{T}$, the decision maker cannot access the optimal solution $\boldsymbol{y}^{*}$ for the current period.
+In this case, we will update the approximations $\boldsymbol{u}^{t}$ and $\boldsymbol{d}^{t}$ by simple first-order computations in a “greedy” fashion as described in the algorithm: Once a type-$j$ customer arrives at period $t-1$, we subtract the number of future type-$j$ arrivals by one, i.e., $\boldsymbol{d}^{t}=\boldsymbol{d}^{t-1}-\boldsymbol{e}_{j}$; Once a type-$j$ customer is accepted at period $t-1$, we subtract the number of future accepted type-$j$ arrivals by one, i.e., $\boldsymbol{u}^{t}=\boldsymbol{u}^{t-1}-\boldsymbol{e}_{j}$.
+Given the approximations $\boldsymbol{u}^{t}$ and $\boldsymbol{d}^{t}$, we adopt the idea of the “argmax” policy in Arlotto and Gurvich (2019) and Vera et al.
+(2021) to make accept/reject decisions.
+At each period $t$, the decision maker observes the arrival type $j$ and accepts the request only if it is feasible to do so (i.e., $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$) and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$.
+Note that $\boldsymbol{u}_{t}$ approximates the number of customers that should be accepted and $\boldsymbol{d}_{t}$ approximates the future demand.
+Intuitively, $d_{j}^{t}-u_{j}^{t}$ represents the number of type $j$ requests that should be rejected and the decision maker accepts the request $j$ if and only if more should be accepted than rejected (i.e., $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$).
 In the following, we use $\mathcal{A}$ to denote the AIR policy in Algorithm 1.
@@
 (2014) and Gupta and Molinaro (2014).
-Since empirical estimations fluctuate drastically at the beginning of the time horizon, the learning time set is designed to update the estimations promptly to avoid the accumulation of error.
+Since empirical estimations fluctuate drastically at the beginning of the time horizon, the learning time set is designed to update the estimations promptly to avoid the accumulation of learning error.
 In contrast to the learning time set, the approximation time set $\mathcal{T}_{A}$ is concentrated at the end of the time horizon.
 It shares some similarities with the NRM literature, e.g., Jasin and Kumar (2012) and Bumpensanti and Wang (2020).
-Since the solution to $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ usually changes drastically when the inventory is scarce, it is crucial to have a good approximation of the optimal solution at the end of the time horizon.
-The approximation time set is thus designed to control the deviation of the approximated solution from the true optimal fluid solution.
+At the end of the time horizon, the solution to $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}}^{t})$ usually changes drastically because of the scarce inventory.
+Since the first-order computations may incur significant errors in this case, the approximation time set is thus designed to control the deviation of the approximated solution from the true optimal fluid solution.
 The total number of resolving is $|\mathcal{T}|=K_{L}+K_{A}+1=\mathcal{O}(\log\log T)$.
@@
 Given the resolving schedule $\mathcal{T}$ with $\alpha\in(0,1)$ and $\beta\in(\frac{1}{2},1)$, the regret of the AIR policy is $\mathcal{O}(1)$.
-According to Theorem 1, the AIR algorithm can achieve a constant regret by solving $\mathcal{O}(\log\log T)$ LPs. In the following, we present the proof idea of Theorem 1.
+According to Theorem 1, the AIR algorithm can achieve a constant regret by solving $\mathcal{O}(\log\log T)$ LPs.
+Before proceeding, we provide an overview of the proof idea of Theorem 1 in four steps.
+First, we decompose the regret by periods, and identify events under which a revenue loss occurs.
+By doing so, we reduce the analysis to bounding the probability of such events.
+Second, we prove that $\boldsymbol{d}^{t}$ in Algorithm 1 is a good approximation of future arrival numbers.
+Third, we identify that $\boldsymbol{u}^{t}$ is an optimal solution of a surrogate LP with high probability, and hence can show that $\boldsymbol{u}^{t}$ is a good approximation of the optimal solution of the fluid problem.
+Lastly, given that $\boldsymbol{u}^{t}$ and $\boldsymbol{d}^{t}$ are good approximations, we can prove that the event probability in the first step is small under the argmax decision, resulting in a constant regret.
+In the following subsection, we present a more detailed proof of Theorem 1 step by step.
+Before that, we first provide remarks to compare our techniques with the literature reviewed in Section 1.1 to highlight our contribution.
+Remark 1 (Comparison with Literature).
+First, Bumpensanti and Wang (2020) consider the problem where the probabilities are known at the beginning, and provide a constant-regret algorithm with an $\mathcal{O}(\log\log T)$ resolving schedule similar to our approximation time set.
+They use the thresholding technique to adjust the acceptance probabilities in the PAC control: When the probability is lower (higher, resp.) than a threshold, the probability will be adjusted to 0 (1, resp.).
+The high-level philosophy of the thresholding policy is similar to the argmax policy, which can be seen as a thresholding action with $0.5$ as the threshold.
+However, different from our policy, their policy between two resolving periods is static (i.e., the acceptance probabilities are not updated), and the induction proof in Bumpensanti and Wang (2020) highly relies on the knowledge of arrival probabilities.
+Even if we complement the resolving time set with our learning time set, their proof cannot easily be adapted to the unknown-probability case.
+Indeed, in Bumpensanti and Wang (2020), they propose a way to deal with the unknown-probability case by requiring per-period resolving rather than infrequent resolving.
+Second, for the known-probability case, Banerjee and Freund (2024) provide a lazy-resolving algorithm achieving a constant regret with $\mathcal{O}(\log\log T)$ expected number of resolvings.
+Their algorithm sets lower confidence bounds for the numbers of accepted customers of different types based on the optimal solution of the fluid model, and updates the bounds when the number of accepted customers exceeds the corresponding bound.
+The proof cannot be extended to the unknown-probability case because the lower confidence bounds is sensitive to the estimation error of the arrival probabilities.
+Third, for the unknown-probability case, Xie et al.
+(2023) achieve constant regret by implementing the argmax policy with per-period resolving.
+However, the proof cannot be directly used for our infrequent resolving schedule because the empirical estimation is infrequently updated and hence the optimal solution to the fluid model is not available in most periods.
+To deal with this problem, we provide a novel idea to identify a surrogate LP which admits the approximated solution as an optimal solution with high probability, such that the approximated solution can be compared with the optimal solution of the fluid model.
+∎
 2.3 Proof of Regret Bound
@@
 According to Proposition 2, the demand estimation error is relatively small with high probability.
-Both bounds consist of two parts: The first term is due to the gap between the empirical probability $\hat{\boldsymbol{p}}^{t}$ and the underlying true probability $\boldsymbol{p}$; the second term is due to the deviation of the random future demand $\boldsymbol{Z}^{t}$ from the expected demand $(T-t+1)\hat{\boldsymbol{p}}^{t}$.
+To explain the bounds in Proposition 2, we note that each bound consists of two terms: The first term is due to the gap between the empirically estimated probability and the underlying true probability; the second term is due to the deviation of the random future demand from the empirically estimated future demand.
+Note that since the empirical probability is also infrequently updated, the demand estimation $\boldsymbol{d}^{t}$ contains $\hat{\boldsymbol{p}}^{T_{k-1}}$ rather than $\hat{\boldsymbol{p}}^{t}$, and hence the concentration bounds in Proposition 2 contains $T_{k-1}$.
 Step 3: Surrogate LP for $\boldsymbol{u}^{t}$.
-In the literature on argmax policies (e.g., Vera and Banerjee 2021 and Vera et al.
+In the literature on argmax policies (e.g., Vera and Banerjee 2020 and Vera et al.
 2021), the proposed policy makes decisions based on the optimal solution of the fluid problem at each period.
@@
     $\boldsymbol{u}^{t}$ is an optimal solution of the LP $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$.
-As Proposition 3 shows, the approximated solution $\boldsymbol{u}^{t}$ is optimal to a surrogate LP, $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$, with high probability.
+The first part in Proposition 3 is proved by concentration equalities.
+The second part is proved by induction: At the resolving period $T_{k-1}$, $\boldsymbol{u}^{T_{k-1}}$ is certainly the optimal solution of $\phi(\boldsymbol{b}^{T_{k-1}},\boldsymbol{d}^{T_{k-1}})$.
+Then, due to the argmax policy and the designed first-order computations in Algorithm 1, we can show that the statement holds for any period $\ell\in(T_{k-1},t]$ as long as $\boldsymbol{d}^{t}\geq\boldsymbol{2}$.
+Proposition 3 shows, the approximated solution $\boldsymbol{u}^{t}$ is optimal to a surrogate LP, $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$, with high probability.
 Recall that we prove that the demand estimation $\boldsymbol{d}^{t}$ is not far away from the random demand $\boldsymbol{Z}^{t}$.
@@
     $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$.
+The proof idea of Proposition 4 is as follows: Consider a type-$j$ customer is accepted at period $t$ and the good event happens.
+In this case, we have $\boldsymbol{d}^{t}\approx\boldsymbol{Z}^{t}$, $\boldsymbol{d}^{t}\geq\frac{T-t+1}{2}\boldsymbol{p}$ and that $\boldsymbol{u}^{t}$ is an optimal solution of $\phi(\boldsymbol{b}^{t},\boldsymbol{d}^{t})$.
+Since $\boldsymbol{d}^{t}\approx\boldsymbol{Z}^{t}$, we have $\boldsymbol{u}^{t}\approx\tilde{\boldsymbol{y}}^{*}$, where $\tilde{\boldsymbol{y}}^{*}$ denotes the optimal solution to $\phi(\boldsymbol{b}^{t},\boldsymbol{Z}^{t})$.
+Then, when $T-t+1$ is greater than a constant, we have $\tilde{y}^{*}_{j}\approx u_{j}^{t}\geq d_{j}^{t}/2\geq\frac{T-t+1}{4}p_{j}\geq 1$, resulting in the results in Proposition 4.
 According to Proposition 4, for most periods, under the good event, we have $\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=1$ and $Z_{j^{t}}^{t}-\mathfrak{S}(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t})\geq 1$ if $x_{j^{t}}^{t}=0$, resulting in $\Delta(\boldsymbol{b}^{t},\boldsymbol{Z}^{t},j^{t},x_{j^{t}}^{t})=0$.
@@
 and
-$\mathcal{T}^{F}_{A}(M)=\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{M-2}}\right\rceil\right\}.$
+$\mathcal{T}^{F}_{A}(M)=\left\{\left\lceil T-T^{\beta}\right\rceil,\left\lceil T-T^{\beta^{2}}\right\rceil,\dots,\left\lceil T-T^{\beta^{M-2}}\right\rceil\right\},$
+where $\epsilon$ can be any positive constant.
 To facilitate understanding, we illustrate the resolving schedule in Figure 2.
@@
 Then, at time point $\lceil T/2\rceil$, we solve the fluid model with a more accurate estimation based on the collected samples.
+The insight behind this choice is that, at early stages, there is plenty of inventory and wrong actions can be made up by the remaining periods.
+Specifically, if we accept too much type-$j$ customers at early periods, then we can reject more in the remaining periods to make up.
+However, at late periods, we cannot tolerate such many wrong actions.
 As for the approximation set, in order to correct the approximation error in time, the smaller time points (further away from $T$) are more crucial.
@@
 Theorem 2.
-For $M\geq 2$, given the finite-resolving schedule $\mathcal{T}^{F}(M)$ with $\beta\in(\frac{1}{2},1)$ the regret of the AIR policy is $\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-2}})$.
+For $M\geq 2$ and any $\epsilon>0$, given the finite-resolving schedule $\mathcal{T}^{F}(M)$ with $\beta\in(\frac{1}{2},1)$ the regret of the AIR policy is $\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-2}})$.
 The proof of Theorem 2 is similar to Theorem 1 and is given in Appendix B.7.
@@
 Therefore, Theorem 2 indicates that the AIR policy can achieve outstanding performance with very limited resolvings.
+Remark 2 (Role of Finite-Support Assumption).
+As mentioned in the introduction, we assume that the underlying distribution $\mathcal{P}$ has finite supports.
+We would like to explain the role of the assumption in the above proof.
+First, the finite-support assumption enables us to aggregate customers by types and derive a tractable fluid model (4) with finite variables.
+If the distribution $\mathcal{P}$ has infinite supports, then the fluid model becomes an intractable infinite linear program.
+In that case, the typical approach is to work on the dual problem instead of the primal (see, Li and Ye 2022, Bray 2024 and Ma et al.
+2024).
+However, in the degenerate case, the linkage between the optimal primal solution and the optimal dual solution becomes relatively weak, resulting in large regrets of the above algorithms.
+Second, the finite-support assumption ensures that the arrival probability of each customer type is positive, which plays an important role in Proposition 4.
+As we explained after Proposition 4, one of the key parts is to deduce $\frac{T-t+1}{4}p_{j}\geq 1$ when $T-t+1$ is greater than a constant independent of $T$.
+However, for the continuous-support case, the probability measure of each type is zero and hence the proof cannot hold.
+Indeed, as Bray (2024) showed, under the continuous-support assumption, the regret of any policy is lower bounded by $\Omega(\log T)$, i.e., a constant-regret algorithm is impossible. ∎
 3 Known Arrival Probabilities
@@
      Observe arrival type $j$ and set $\boldsymbol{N}^{t+1}\leftarrow\boldsymbol{N}^{t}+\boldsymbol{e}_{j}$.
-     if $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$ and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
+     if  $\boldsymbol{A}_{j}\leq\boldsymbol{b}^{t}$, $u_{j}^{t}>1$, and $u_{j}^{t}\geq d_{j}^{t}-u_{j}^{t}$ then $\triangleright$ Argmax between $u_{j}^{t}$ and $d_{j}^{t}-u_{j}^{t}$
          Accept the request.
@@
 Theorem 3 (Regret Bound for NRM).
-Given the arrival probabilities and the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(\frac{1}{2},1)$, the regret of the AIR-KP policy is $\mathcal{O}(1)$.
+For any $\epsilon>0$, given the arrival probabilities and the resolving schedule $\mathcal{T}^{\mathcal{K}}$ with $\beta\in(\frac{1}{2},1)$, the regret of the AIR-KP policy is $\mathcal{O}(1)$.
 The idea of the proof of Theorem 3 is similar to that of Theorem 1, which is described in Section 2.3.
@@
 However, different from Jasin and Kumar (2012), our resolving schedule also works for the degenerate case.
+Moreover, in the following proposition, we provide a lower bound for the regret of the AIR-KP policy with any two resolving time points.
+Proposition 5.
+Given any resolving schedule with no more than two resolvings, the regret of the AIR-KP policy is $\Omega(T^{1/4})$.
+According to Theorem 4 and Proposition 5, the regret bound $\mathcal{O}(T^{(1/2+\epsilon)^{M}})$ (with $\epsilon\to 0$) is nearly tight when $M=2$.
 3.2 Discussion on Resolving Schedules
@@
 In the following, we compare the performance of the above algorithms.
+[^4] Note that the distribution assumptions and the non-degeneracy assumption in the Gao et al.
+(2024) and Ma et al.
+(2024) may not be satisfied in the numerical experiments.
 For each parameter set, we run 2,000 simulations.
@@
 First, in Figure 4, we fix the time horizon $T$ as $50,000$, and test the algorithms as the budget factor $\rho$ changes.
-According to Figure 4, the regret of most policies is large when $\rho$ is close to 0.5.
-Due to the setup of the single-resource problem, the fluid problem at time $1$ is degenerate when $\rho=0.5$.
+Due to the setup of the single-resource problem, the fluid problem at time $1$ is degenerate when $\rho$ takes the value $0.5$ and hence the non-degeneracy measure $\delta$ equals $|\rho-0.5|$.
+According to Figure 4, the regret of most policies is large when $\rho$ is close to 0.5 (i.e, $\delta\to 0$).
 Therefore, the degenerate case is the hardest case for most policies.
@@
 Table 4: Resolving Schedule $\mathcal{T}$ with $\alpha=\beta=0.7$.
+Then, we investigate the impact of the hyper-parameters (i.e., $\alpha$ and $\beta$) on the regret in Table 5.
+According to Table 5, as $\alpha$ or $\beta$ increases, the regret first decreases steeply and then gradually becomes steady.
+Therefore, to implement our policy in practice, it is recommended to set slightly large hyper-parameters ($\alpha$ and $\beta$).
+$\alpha$ Regret # of Resolvings
+0.15 887.4 10
+0.25 211.7 10
+0.35 79.6 11
+0.45 13.6 11
+0.55 2.4 12
+0.65 2.2 14
+0.75 2.0 16
+0.85 2.0 22
+0.95 1.8 45
+(a) Fix $\beta=0.7$
+$\beta$ Regret # of Resolvings
+0.55 9.2 12
+0.60 5.6 13
+0.65 3.3 14
+0.70 2.1 15
+0.75 1.9 16
+0.80 1.8 18
+0.85 1.8 22
+0.90 1.7 28
+0.95 1.7 45
+(b) Fix $\alpha=0.7$
+Table 5: Regret of AIR Policy as $\alpha$ or $\beta$ Changes ($T=30,000$).
+[…]
+Figure 5: Regret under Different Policies as Functions of $T$ when $m=10$, $n=50$ and $\alpha=\beta=0.9$.
 Lastly, in order to identify the impact of $n$, we consider the case when $m=10$ and $n=50$, and compare the AIR policy with LP-free policies.
@@
 The reason is that as the number of customer types increases, the optimal solution of the fluid problem $\phi(\boldsymbol{b}^{t},(T-t+1)\hat{\boldsymbol{p}})$ changes more significantly and frequently and hence the resolving frequency should be slightly increased.
-[…]
-Figure 5: Regret under Different Policies as Functions of $T$ when $m=10$, $n=50$ and $\alpha=\beta=0.9$.
 4.2 Finite Resolving
@@
   1. 1.
-    For the time interval $\left[1,\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right)$, Algorithm 1 incurs at most $T^{(1/2+\epsilon)\cdot\beta^{M-2}}\cdot r_{\phi}$ regret.
+    For the time interval $\left[1,\left\lceil T^{(1/2+\epsilon)\cdot\beta^{M-2}}\right\rceil\right)$, Algorithm 1 rejects all customers and incurs at most $T^{(1/2+\epsilon)\cdot\beta^{M-2}}\cdot r_{\phi}$ regret.
   2. 2.
@@
 $\leq\frac{c_{8}\pi^{2}}{6}r_{\phi}+T^{(1/2+\epsilon)\cdot\beta^{M-1}}r_{\phi}=\mathcal{O}(T^{(1/2+\epsilon)\cdot\beta^{M-1}}).$
-B.13 Proof of Lemma 3
+B.13 Proof of Proposition 5
+According to Theorem 3, we have $\mathbb{E}\left[\phi(T\boldsymbol{\rho},\boldsymbol{Z}^{1})\right]-V^{*}(T)=O(1)$, i.e., the difference between the hindsight problem and the optimal problem is upper bounded by a constant.
+Thus, in the following, we will compare the policy with the hindsight problem.
+Let $T_{1}$ and $T_{2}$ (with $T_{1}<T_{2}$) denote the two resolving time points.
+Similar to Bumpensanti and Wang (2020), in this proof, we consider two types of customers and one type of resource.
+The arrival probabilities of either customer type is $0.5$.
+The rewards of the two types of customers are $r_{1}=2$ and $r_{2}=1$, and each customer consumes one unit of resource.
+The budget per period is $0.5$, resulting in a degenerate case.
+We start with the case where $T_{1}=\omega(T^{1/4})$, which means that $\lim_{T\to\infty}\frac{T_{1}}{T^{1/4}}=\infty$.
+In this case, due to the initialization in Algorithm 2, the algorithm will reject all customers during the time interval $[1,T_{1})$.
+Let $\Lambda_{j}(t_{1},t_{2})$ denote the random number of type-$j$ arrivals during the time interval $[t_{1},t_{2}]$.
+We consider the event where $\frac{1}{2}(T_{1}-1)-\sqrt{T_{1}-1}\leq\Lambda_{1}(1,T_{1}-1)\leq\frac{1}{2}(T_{1}-1)$ and $\Lambda_{1}(T_{1},T)\leq\frac{1}{2}(T-T_{1}+1)$.
+In this case, we have $\Lambda_{1}(1,T)=\Lambda_{1}(1,T_{1}-1)+\Lambda_{1}(T_{1},T)\leq\frac{1}{2}T$ and hence the hindsight problem will accept all type-1 customers.
+However, the algorithm rejects $\Lambda_{1}(1,T_{1}-1)$ type-1 customers, and hence the corresponding revenue loss is at least $\Lambda_{1}(1,T_{1}-1)\geq\frac{1}{2}(T_{1}-1)-\sqrt{T_{1}-1}=\Omega(T^{1/4})$.
+Then, similar to Bumpensanti and Wang (2020), according to the Berry-Esseen theorem (see Shevtsova 2011), we bound the probability of the event as follows:
+$\mathbb{P}\left(\frac{1}{2}(T_{1}-1)-\sqrt{T_{1}-1}\leq\Lambda_{1}(1,T_{1}-1)\leq\frac{1}{2}(T_{1}-1)\;\&\;\Lambda_{1}(T_{1},T)\leq\frac{1}{2}(T-T_{1}+1)\right)$
+$=$ $\mathbb{P}\left(-\frac{\sqrt{T_{1}-1}}{\frac{1}{2}\sqrt{T_{1}-1}}\leq\frac{\Lambda_{1}(1,T_{1}-1)-\frac{1}{2}(T_{1}-1)}{\frac{1}{2}\sqrt{T_{1}-1}}\leq 0\right)\cdot\mathbb{P}\left(\frac{\Lambda_{1}(T_{1},T)-\frac{1}{2}(T-T_{1}+1)}{\frac{1}{2}\sqrt{T-T_{1}+1}}\leq 0\right)$
+$\geq$ $\left(0.477-2\times\frac{0.4748}{\frac{1}{8}\sqrt{T_{1}-1}}\right)\cdot\left(0.5-\frac{0.4748}{\frac{1}{8}\sqrt{T-T_{1}+1}}\right),$
+which is greater than a constant when $T$ is no less than a threshold.
+Therefore, the regret will be the order of $\Omega(T^{1/4})$.
+Then, we consider the case where $T_{1}=O(T^{1/4})$, and study different choices of $T_{2}$.
+Case I: There exists a constant $c_{12}<1$ such that $T_{1}\leq T_{2}\leq c_{12}\cdot T$.
+In this case, we consider the event when $\frac{1}{2}(T_{2}-1)\leq\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)+\sqrt{T_{2}-1}$ and $\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)-\sqrt{T}$.
+Under this event, the total number of type-1 arrivals is $\Lambda_{1}(1,T_{2}-1)+\Lambda_{1}(T_{2},T)\leq\frac{1}{2}T+\sqrt{T_{2}-1}-\sqrt{T}\leq\frac{1}{2}T-(1-\sqrt{c_{12}})\sqrt{T}$, and hence the hindsight optimal policy accepts at least $(1-\sqrt{c_{12}})\sqrt{T}$ type-2 customers.
+Then, we analyze the performance of the AIR-KP policy with the resolving schedule $\{T_{1},T_{2}\}$.
+During the time interval $[1,T_{1})$, the policy rejects all customers.
+Then, the algorithm resolves the fluid model and derives the solution $\boldsymbol{u}^{T_{1}}=\left(\frac{1}{2}(T-T_{1}),\frac{1}{2}T_{1}\right)$ and the demand estimation $\boldsymbol{d}^{T_{1}}=\left(\frac{1}{2}(T-T_{1}),\frac{1}{2}(T-T_{1})\right)$.
+Thus, during the interval $[T_{1},T_{2})$, the algorithm will accept all type-1 customers and no greater than $\frac{1}{2}T_{1}$ type-2 customers, resulting in $\nu\geq\Lambda_{1}(1,T_{2}-1)-T_{1}$ resource consumption.
+When the algorithm resolves the fluid model at time $T_{2}$, we can derive the solution $\boldsymbol{u}^{T_{2}}$ with $u_{2}^{T_{2}}=\max\left\{\frac{1}{2}T-\nu-\frac{1}{2}(T-T_{2}+1),0\right\}\leq\max\left\{\frac{1}{2}(T_{2}-1)-\frac{1}{2}(T_{2}-1)-T_{1},0\right\}\leq T_{1}$.
+Therefore, the algorithm will accept at most $\frac{3}{2}T_{1}$ type-2 customers and induce at least $(1-\sqrt{c_{12}})\sqrt{T}-\frac{3}{2}T_{1}=\Omega(\sqrt{T})$ revenue loss under this event.
+Then, we provide a lower bound for the event probability as follows (Let $\Phi(\cdot)$ denote the c.d.f of a standard normal distribution):
+$\mathbb{P}\left(\frac{1}{2}(T_{2}-1)\leq\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)+\sqrt{T_{2}-1}\;\&\;\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)-\sqrt{T}\right)$
+$=$ $\mathbb{P}\left(0\leq\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq 2\right)\cdot\mathbb{P}\left(\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq-\frac{\sqrt{T}}{\frac{1}{2}\sqrt{T-T_{2}+1}}\right)$
+$\geq$ $\mathbb{P}\left(0\leq\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq 2\right)\cdot\mathbb{P}\left(\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq-\frac{2}{\sqrt{1-c_{12}}}\right)$
+$\geq$ $\left(0.477-2\times\frac{0.4748}{\frac{1}{8}\sqrt{T_{2}-1}}\right)\cdot\left(\Phi\left(-\frac{2}{\sqrt{1-c_{12}}}\right)-\frac{0.4748}{\frac{1}{8}\sqrt{T-T_{2}+1}}\right),$
+which is greater than a constant when $T$ is above a threshold. Therefore, the regret of the algorithm is at least $\Omega(\sqrt{T})$.
+Case II: $T-T_{2}+1=\Omega(\sqrt{T})$ and $T-T_{2}+1=o(T)$.
+In this case, we consider the event when $\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T-T_{2}+1}$ and $\frac{1}{2}(T-T_{2}+1)+\frac{1}{2}\sqrt{T-T_{2}+1}\leq\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)+\sqrt{T-T_{2}+1}$.
+Under this event, we have $\Lambda_{1}(1,T)=\Lambda_{1}(1,T_{2}-1)+\Lambda_{1}(T_{2},T)\leq\frac{1}{2}T$, and hence the hindsight optimal policy accepts all type-1 customers.
+Then, we analyze the performance of the AIR-KP policy with the resolving schedule $\{T_{1},T_{2}\}$.
+Similar to Case I, the algorithm rejects all customers during the interval $[1,T_{1})$.
+During the interval $[T_{1},T_{2})$, the algorithm accepts all type-1 customers and at most $\frac{1}{2}T_{1}$ type-2 customers.
+Then, the algorithm resolves the fluid model and get the solution $\boldsymbol{u}^{T_{2}}$ with $u_{1}^{T_{2}}=\min\left\{\frac{1}{2}T-\nu,\frac{1}{2}(T-T_{2}+1)\right\}\leq\frac{1}{2}(T-T_{2}+1)$.
+However, since $\Lambda_{1}(T_{2},T)\geq\frac{1}{2}(T-T_{2}+1)+\frac{1}{2}\sqrt{T-T_{2}+1}$, the algorithm rejects at least $\frac{1}{2}\sqrt{T-T_{2}+1}$ type-1 customers, resulting in $\Omega(T^{1/4})$ revenue loss.
+Then, given a sufficiently large $T$, we provide lower bound for the event probability as follows:
+$\mathbb{P}\left(\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T-T_{2}+1}\;\&\;\frac{1}{2}(T-T_{2}+1)+\frac{1}{2}\sqrt{T-T_{2}+1}\leq\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T-T_{2}+1)+\sqrt{T-T_{2}+1}\right)$
+$=$ $\mathbb{P}\left(\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq-\frac{\sqrt{T-T_{2}+1}}{\frac{1}{2}\sqrt{T_{2}-1}}\right)\cdot\mathbb{P}\left(1\leq\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq 2\right)$
+$\geq$ $\mathbb{P}\left(\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq-2\right)\cdot\mathbb{P}\left(1\leq\frac{\Lambda_{1}(T_{2},T)-\frac{1}{2}(T-T_{2}+1)}{\frac{1}{2}\sqrt{T-T_{2}+1}}\leq 2\right)$
+$\geq$ $\left(0.022-\frac{0.4748}{\frac{1}{8}\sqrt{T_{2}-1}}\right)\cdot\left(0.135-2\times\frac{0.4748}{\frac{1}{8}\sqrt{T-T_{2}+1}}\right),$
+which is greater than a constant when $T$ is above a threshold. Therefore, the regret of the algorithm is at least $\Omega(T^{1/4})$.
+Case III: $T-T_{2}+1=o(\sqrt{T})$.
+In this case, we consider the event when $\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T_{2}-1}$.
+Under this event, given a sufficiently large $T$, we have $\Lambda_{1}(1,T)=\Lambda_{1}(1,T_{2}-1)+\Lambda_{1}(T_{2},T)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T_{2}-1}+(T-T_{2}+1)\leq\frac{1}{2}(T_{2}-1)-\frac{1}{2}\sqrt{T_{2}-1}$, and hence the hindsight optimal policy accepts at least $\frac{1}{2}\sqrt{T_{2}-1}$ type-2 customers.
+Then, similar to the analysis in Case I, the algorithm rejects all customers during the interval $[1,T_{1})$.
+During the interval $[T_{1},T_{2})$, the algorithm accepts all type-1 customers and at most $\frac{1}{2}T_{1}$ type-2 customers.
+During the interval $[T_{2},T]$, the algorithm accepts at most $T-T_{2}+1$ type-2 customers.
+Thus, the algorithm accepts at most $\frac{1}{2}T_{1}+T-T_{2}+1$ type-2 customers, resulting in $\frac{1}{2}\sqrt{T_{2}-1}-\frac{1}{2}T_{1}-(T-T_{2}+1)=\Omega(\sqrt{T})-\mathcal{O}(T^{1/4})-o(T^{1/4})=\Omega(\sqrt{T})$.
+Then, we provide a lower bound for the event probability as follows:
+$\mathbb{P}\left(\Lambda_{1}(1,T_{2}-1)\leq\frac{1}{2}(T_{2}-1)-\sqrt{T_{2}-1}\right)=\mathbb{P}\left(\frac{\Lambda_{1}(1,T_{2}-1)-\frac{1}{2}(T_{2}-1)}{\frac{1}{2}\sqrt{T_{2}-1}}\leq-2\right)\geq\left(0.022-\frac{0.4748}{\frac{1}{8}\sqrt{T_{2}-1}}\right),$
+which is greater than a constant when $T$ is greater than a threshold. Therefore, the regret will be $\Omega(\sqrt{T})$.
+To summarize, given two resolving time points, the regret of the AIR-KP policy is $\Omega(T^{1/4})$. ∎
+B.14 Proof of Lemma 3
 For the known-probability problem, we need to prove similar results in Lemmas 7 and 8. Similar to Lemma 7, we have
@@
 ∎
-B.14 Proof of Proposition 5
+B.15 Proof of Proposition 6
 For the unknown-probability case, we only need to prove similar results in Propositions 3 and 4.
@@
 Then, similar to Proposition 4, we can finish the proof.
-B.15 Proof of Lemma 4
+B.16 Proof of Lemma 4
 Similar to Lemma 3, we only need the following inequalities:
@@
 Then, following a similar proof, we can derive the constant bound $\mathcal{O}(1)$. ∎
-B.16 Proof of Proposition 6
+B.17 Proof of Proposition 7
 Similar to Proposition 6, we only need the following inequalities: